### v3

# Adaptive Transfer Clustering: A Unified Framework

Abstract
We propose a general transfer learning framework for clustering given a main dataset and an auxiliary one about the same subjects. The two datasets may reflect similar but different latent grouping structures of the subjects. We propose an adaptive transfer clustering (ATC) algorithm that automatically leverages the commonality in the presence of unknown discrepancy, by optimizing an estimated bias-variance decomposition. It applies to a broad class of statistical models including Gaussian mixture models, stochastic block models, and latent class models. A theoretical analysis proves the optimality of ATC under the Gaussian mixture model and explicitly quantifies the benefit of transfer. Extensive simulations and real data experiments confirm our method’s effectiveness in various scenarios.
Keywords: Multiview clustering; Transfer learning; Adaptation; Bootstrap.
1 Introduction
In recent years, data collection from multiple sources or views, each offering unique insights into the underlying structure, has become increasingly common. As a result, transfer learning has gained prominence as a powerful framework in machine learning [49]. While transfer learning has been widely applied to supervised settings [53] and semi-supervised settings [9, 40, 70], its application to unsupervised settings still remains at the infant stage. In this background, clustering has emerged as a central unsupervised task with applications in medical imaging [50, 32, 63], genetics [18, 45, 60, 54], sentiment analysis [68, 12], to name just a few.
Existing transfer learning approaches for clustering primarily focus on the feature space — where the target domain and source domain are different but related through shared or similar parameters. Gaussian Mixture Model (GMM) [51] is one of the most canonical models for transfer clustering, which itself plays a fundamental role in statistics. A line of works [59, 56] on GMMs apply EM-type algorithms that leverage similar mean or covariance structures in source domain to improve the clustering performance of target data. Specifically, [56] provides theoretical guarantees by assuming closeness of discriminant coefficients in Euclidean distance.
Nonetheless, these works typically assume similar parameter structures in the source and the target domains, which contain different subjects. Little attention has been given to the cases where the target and the source data consist of different aspects of the same subjects, which have similar but different labels and might even follow different models. Such problems naturally arise in many practical settings. For instance, in social science we may be interested in clustering a group of people based on their friendship network and attributes (e.g., age, education, occupation). However, the network and attributes often exhibit different signal strengths and may not reflect the same community structure. Another example in neuroimaging, as noted by [50], is that the community structures within brain networks may differ between subjects due to their physiological differences, responses to changing environmental conditions, and variations in imaging instruments. A related work [10] investigates the fundamental limits of community detection in multi-layer networks under label flipping for individual networks. However, their method is not adaptive to an unknown flipping probability. There is a crucial need for procedures that can adaptively leverage useful information from the source domain to assist clustering in the target domain.
Main contributions
Motivated by the above need, we propose a general framework for adaptive transfer clustering. Our target data $\mathbf{X}_{0}$ and source data $\mathbf{X}_{1}$ reflect different features of the same set of $n$ subjects, where each dataset has $K$ latent groups. In this framework, $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$ may come from different mixture models, where a discrepancy parameter $\varepsilon$ indicates the proportion of mismatched latent labels $\mathbf{Z}^{*}_{0},\mathbf{Z}^{*}_{1}\in[K]^{n}$. Our goal is to estimate $\mathbf{Z}_{0}^{*}$ from $(\mathbf{X}_{0},\mathbf{X}_{1})$. The key challenge is how to adaptively leverage the information in $\mathbf{X}_{1}$ to cluster $\mathbf{X}_{0}$, without knowing the discrepancy $\varepsilon$. Intuitively, when $\varepsilon=0$, i.e., the labels are perfectly matched, we should “pool” the data together; when $\varepsilon$ is large enough, i.e., little label information about $\mathbf{X}_{0}$ can be inferred from $\mathbf{X}_{1}$, we may discard $\mathbf{X}_{1}$ and perform clustering on the target data $\mathbf{X}_{0}$ only. Inspired by this, we develop a model-based method with a penalty term encouraging the similarity of $\mathbf{Z}^{*}_{0}$ and $\mathbf{Z}^{*}_{1}$. Informally, we try to minimize the following objective with $\lambda>0$:
log posterier of ($-\text{log posterier of }(\mathbf{Z}_{0}\mid\mathbf{X}_{0})-\text{log posterier of }(\mathbf{Z}_{1}\mid\mathbf{X}_{1})+\lambda\cdot\text{penalty}\left(\mathbf{Z}_{0},\mathbf{Z}_{1}\right).$ (1)
To summarize, our contributions are two-fold:
  • Methodology: Our framework is general to incorporate arbitrary mixture distributions of target and source data, making it applicable to a broad class of practical models including Gaussian mixture models, latent class models, and contextual stochastic block models. We design an adaptive procedure, named Adaptive Transfer Clustering or ATC for short, that can select the crucial parameter $\lambda$ in (1) agonistic to the level of discrepancy $\varepsilon$. The key ingredient of achieving adaptivity is a combination of Goldenshluger-Lepski method [24] and parametric bootstrap.
  • Theory: We establish a sharp clustering error rate for ATC in cases where both the target and source data are generated from two-component $d$-dimensional GMMs. Denote by SNR the signal-to-noise ratio such that the best estimator using only the target data achieves a rate of $\exp\left(-\textsf{SNR}\left(1+o(1)\right)\right)$ as $\textsf{SNR}\rightarrow\infty$ [42, 44, 21]. We show that the optimal rate in the transfer learning setting is
    $\exp\left(-\textsf{SNR}\min\left\{1+\frac{\log\left(1/\varepsilon\right)}{4\textsf{SNR}},2\right\}\left(1+o(1)\right)\right),$
    and it can be achieved by our ATC procedure without knowing $\varepsilon$. This rate is always better than the aforementioned target-only rate.
Related work
Our work is broadly related to the emerging area of unsupervised learning in the transfer learning setting [52, 45, 15, 30]. Below, we outline a few related topics with a non-exhaustive review of relevant literature.
  • Multi-view clustering: Our work connects to the paradigm of multi-view clustering, which aims to cluster multiple datasets representing different aspects (views) of the same subjects [4]. Existing works provide discriminative approaches that optimize carefully designed objective functions to maximize intra-cluster similarity and minimize inter-cluster similarity [37, 58], or generative approaches based on EM-type algorithms or Bayesian framework [41, 35, 23, 22]. See [8] for a survey. We will discuss comparisons in detail in Section 3.
  • Network community detection with side information: Our work is related to network community detection literature involving multi-layer networks [47, 10, 29, 11, 33, 25] and contextual networks [64, 5, 48, 69, 61, 46, 3, 62, 31, 6, 43, 14]. In the multi-layer network setting, recent works such as [13, 11] propose multi-view stochastic block models that assume homogeneous label structures across views. [13] uses a variational Bayesian EM algorithm for parameter estimation and clustering. [11] provides algorithmic guarantees for weak recovery and exact recovery for a special case when each network only consists of two clusters. In the contextual network setting, [6] derives an exponential error for Lloyd-type algorithm for contextual stochastic block model (SBM) by assuming same clustering structures. [31] incorporates joint and individual structures in both the network and the covariates, and proposes to use a spectral method followed by a refinement step.
  • Testing for clustering structures: There is also a growing body of theoretical and applied work investigating the testing problems for the common structures across different sources or views [23, 22, 19, 65, 45, 55, 17]. A closely related work is [19], which proposes to test whether two datasets from GMMs share a common clustering structure and derives a sharp detection boundary. However, for the task of clustering, it remains unknown whether a naïve dichotomous strategy, based on their testing procedure, is optimal. See Section 2 for discussion on difference between clustering (label estimation) and testing.
Organization
Section 2 introduces the two-component symmetric univariate Gaussian mixture model as a warm-up example and presents the adaptive transfer clustering algorithm together with theoretical guarantees. Section 3 extends the methodology to a general framework applicable to other models. Section 4 provides theoretical analysis within the general framework, with an application to the two-component symmetric $d$-dimensional Gaussian mixture model fully adaptive to unknown parameters. Section 5 validates the effectiveness of the proposed algorithm through extensive simulations. Section 6 demonstrates the application of the our method to three real-world datasets. Section 7 concludes the paper and discusses future directions. All the proofs of theoretical results are included in the Appendix.
Notation
Throughout the paper, the constants $c_{0},C_{0},c_{1},C_{1},\cdots$ may vary from line to line. We use plain letters $x,z,X,Z,\cdots$ to denote scalers and use boldface ${\bm{x}},{\bm{z}},\mathbf{X},\mathbf{Z},\cdots$ to denote either vectors or matrices. We write $[n]:=\left\{1,2,\cdots,n\right\}$ and $\text{sgn}\left(x\right):=x/\left|x\right|$ for $x\in\mathbb{R}$, or $\text{sgn}\left({\bm{x}}\right):=\left(\text{sgn}\left(x_{1}\right),\cdots,\text{sgn}\left(x_{n}\right)\right)^{\top}$ for ${\bm{x}}\in\mathbb{R}^{n}$. We denote the $\ell_{2}$ vector norm as $\left\|\cdot\right\|:=\left\|\cdot\right\|_{2}$. For nonnegative sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ or $b_{n}\gtrsim a_{n}$ or $a_{n}=O(b_{n})$ if there exists a universal constant $C>0$ such that $a_{n}\leq Cb_{n}$. We write $a_{n}\asymp b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$, and we write $a_{n}=o\left(b_{n}\right)$ or $b_{n}=\omega\left(a_{n}\right)$ if $a_{n}=O\left(c_{n}b_{n}\right)$ for some $c_{n}\rightarrow 0$. In most cases, we omit the subscript of $n$ when it is clear from context.
2 Warm-up: transfer learning for the Gaussian mixture model
In this section, we consider the transfer learning problem for one-dimensional, two-component symmetric Gaussian mixture model as a warm-up example.
2.1 Problem Setup
Let $\{X_{0,i}\}_{i=1}^{n}\subseteq\mathbb{R}$ be i.i.d. samples from a two-component symmetric Gaussian mixture model
$X_{0,i}\sim\frac{1}{2}N(\mu,\sigma^{2})+\frac{1}{2}N(-\mu,\sigma^{2})$
with known parameters $\mu\in\mathbb{R}$ and $\sigma>0$. Assume $\mu\geq 0$ for the sake of identifiability. There exist i.i.d. Rademacher latent variables $\{Z_{0,i}^{*}\}_{i=1}^{n}\subseteq\{\pm 1\}$ such that $X_{0,i}|Z_{0,i}^{*}\sim N(Z_{0,i}^{*}\mu,\sigma^{2})$. The goal of clustering is to recover $\mathbf{Z}_{0}^{*}=(Z_{0,1}^{*},\cdots,Z_{0,n}^{*})$ from $\mathbf{X}_{0}=(X_{0,1},\cdots,X_{0,n})$.
Define the normalized Hamming distance between any pair of label vectors:
$\ell(\mathbf{Z},\mathbf{Z}^{\prime})=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left(Z_{i}\neq Z^{\prime}_{i}\right),\qquad\forall\mathbf{Z},\mathbf{Z}^{\prime}\in\{\pm 1\}^{n}.$
For any estimate ^$\widehat{}\mathbf{Z}$ of $\mathbf{Z}_{0}^{*}$, the quantity (^$\ell(\widehat{}\mathbf{Z},\mathbf{Z}_{0}^{*})$ is often referred to as the misclassification rate. It is easily seen that $\mathrm{sgn}(\mathbf{X}_{0})$ is a maximum a posteriori (MAP) estimate, which achieves the minimum expected misclassification rate $\Phi\left(-\mu/\sigma\right)$, where $\Phi(\cdot)$ is the c.d.f. of standard normal distribution. Define $\textsf{SNR}:=\mu^{2}/\left(2\sigma^{2}\right)$. For our interest, we consider the diverging signal-to-noise ratio regime where $\textsf{SNR}\rightarrow\infty$. On the other hand, we consider $\textsf{SNR}<\log n$ as otherwise $\mathrm{sgn}(\mathbf{X}_{0})$ would lead to exact clustering, a less interesting case. This can be seen from the following relations which holds for $\textsf{SNR}>\log n$ and large $n$:
$\mathbb{E}\ell\left(\mathrm{sgn}(\mathbf{X}_{0}),\mathbf{Z}^{*}\right)=\Phi\left(-\sqrt{2\textsf{SNR}}\right)\leq\exp\left(-\textsf{SNR}\right)<\frac{1}{n}.$
The expected number of misclassified samples is less than $1$, indicating exact clustering with high probability.
Suppose we also collect another batch of i.i.d. samples $\left\{X_{1,i}\right\}_{i=1}^{n}$ from the mixture distribution $\frac{1}{2}N(\mu,\sigma^{2})+\frac{1}{2}N(-\mu,\sigma^{2})$, whose latent variables $\{Z_{1,i}^{*}\}_{i=1}^{n}$ are possibly similar to $\{Z_{0,i}^{*}\}_{i=1}^{n}$ in the following sense:
$\mathbb{P}(Z_{1,i}^{*}\neq Z_{0,i}^{*})\leq\varepsilon.$ (2)
Here $\varepsilon\in[0,1/2]$ is an unknown small number that controls the discrepancy [^1] It worth noting that when $\varepsilon>1/2$, this problem can be reformulated as considering $\left\{X_{0,i}\right\}_{i=1}^{n}$ and $\left\{-X_{1,i}\right\}_{i=1}^{n}$ with discrepancy parameter $\varepsilon^{\prime}=1-\varepsilon$, hence it suffices to restrict ourselves to $\varepsilon\in[0,1/2]$.. We refer to $\mathbf{X}_{0}$ and $\mathbf{X}_{1}=(X_{1,1},\cdots,X_{1,n})$ as the target and the source data, respectively. With the aid of source data, the goal of transfer clustering is to recover $\mathbf{Z}_{0}^{*}$ from $(\mathbf{X}_{0},\mathbf{X}_{1})$. When $\varepsilon=0$, the target data and source data share the same labels, whereas when $\varepsilon=1/2$, the clustering structures between the target data and source data, in the worst scenario, are uninformative. The following are two common estimation strategies which we call independent task learning (ITL) and data pooling (DP), respectively.
Example 2.1 (Independent task learning).
If we only have access to target data $\mathbf{X}_{0}$, the optimal procedure for estimating $\mathbf{Z}^{*}_{0}$ is given by ˇ$\check{}\mathbf{Z}^{\textsf{ITL}}_{0}$, where
$\check{Z}^{\textsf{ITL}}_{0,i}$ $\in\operatorname*{arg\,min}_{u\in\{\pm 1\}}\left\{\frac{(X_{0,i}-u\mu)^{2}}{\sigma^{2}}\right\}=\text{sgn}\left(X_{0,i}\right),\quad\forall i\in[n].$
Example 2.2 (Data pooling).
If we ignore the potential label discrepancy between target data $\mathbf{X}_{0}$ and source data $\mathbf{X}_{1}$, we can estimate $\mathbf{Z}^{*}_{0}$ as if $(\mathbf{X}_{0},\mathbf{X}_{1})$ is from a $2$-dimensional symmetric Gaussian mixture via ˇ$\check{}\mathbf{Z}_{0}^{\textsf{DP}}$, where
$\check{Z}^{\textsf{DP}}_{0,i}$ $\in\operatorname*{arg\,min}_{u\in\{\pm 1\}}\left\{\frac{(X_{0,i}-u\mu)^{2}}{\sigma^{2}}+\frac{(X_{1,i}-u\mu)^{2}}{\sigma^{2}}\right\}=\text{sgn}\left(X_{0,i}+X_{1,i}\right),\quad\forall i\in[n].$
We could obtain the following expected misclassification rate for ˇ$\check{}\mathbf{Z}_{0}^{\textsf{ITL}}$ and ˇ$\check{}\mathbf{Z}_{0}^{\textsf{DP}}$, whose proof is deferred to Appendix C.2.
Proposition 1.
For each $i\in[n]$, we have
$\mathbb{P}\big{(}\check{Z}^{\textsf{ITL}}_{0,i}\neq Z^{*}_{0,i}\big{)}=\Phi\left(-\frac{\mu}{\sigma}\right)\qquad\text{and}\qquad 0\leq\mathbb{P}\big{(}\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i}\big{)}-\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)\leq\frac{\varepsilon}{2}.$
Consequently, we have
$\mathbb{P}\big{(}\check{Z}^{\textsf{ITL}}_{0,i}\neq Z^{*}_{0,i}\big{)}\lesssim\exp\left(-\textsf{SNR}\right)\qquad\text{and}\qquad\mathbb{P}\big{(}\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i}\big{)}\lesssim\exp\left(-2\cdot\textsf{SNR}\right)+\varepsilon.$
A natural question is whether we can do better than ˇ$\check{}\mathbf{Z}_{0}^{\textsf{ITL}}$ and ˇ$\check{}\mathbf{Z}_{0}^{\textsf{DP}}$ in the general case. Next, we propose a transfer clustering procedure to answer this question affirmatively.
2.2 Transfer learning procedure
Consider a special case where $\mathbb{P}(Z_{1,i}^{*}\neq Z_{0,i}^{*}|Z_{0,i}^{*})=\varepsilon$. If $\varepsilon$ is known, then the joint MAP of $(Z_{0,i}^{*},Z_{1,i}^{*})$ is given by
$\operatorname*{arg\,min}_{(u,v)\in\{\pm 1\}^{2}}\left\{\frac{(X_{0,i}-u\mu)^{2}}{2\sigma^{2}}+\frac{(X_{1,i}-v\mu)^{2}}{2\sigma^{2}}+\log\bigg{(}\frac{1-\varepsilon}{\varepsilon}\bigg{)}\mathbb{I}\left(u\neq v\right)\right\}.$ (3)
The term $\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)\mathbb{I}\left(u\neq v\right)$, which involves unknown parameter $\varepsilon$, serves as a penalty term. Motivated by (3), we consider the following procedure to leverage the information from source data. For any $\lambda>0$, define
$\left(\check{Z}^{\lambda}_{0,i},\check{Z}^{\lambda}_{1,i}\right)$ $\in\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{\frac{(X_{0,i}-u\mu)^{2}}{2\sigma^{2}}+\frac{(X_{1,i}-v\mu)^{2}}{2\sigma^{2}}+\lambda\mathbb{I}\left(u\neq v\right)\right\}$
$=\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\mu\left(uX_{0,i}+vX_{1,i}\right)+\lambda\sigma^{2}\mathbb{I}\left(u\neq v\right)\right\},\quad\forall i\in[n]$ (4)
We denote (ˇ$(\check{}\mathbf{Z}^{\lambda}_{0},\check{}\mathbf{Z}^{\lambda}_{1})=\textsf{TC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\lambda,\mu,\sigma\right)$ for the procedure defined by (2.2). Notice that ˇ$\check{}\mathbf{Z}^{\lambda}_{0}$ reduces to the independent task learning estimator when $\lambda=0$ and the data pooling estimator when $\lambda=\infty$, i.e., ˇ$\check{}\mathbf{Z}^{0}_{0}=\check{}\mathbf{Z}^{\textsf{ITL}}_{0}$ and ˇ$\check{}\mathbf{Z}^{\infty}_{0}=\check{}\mathbf{Z}^{\textsf{DP}}_{0}$. As such, ˇ$\check{}\mathbf{Z}^{\lambda}_{0}$ can be viewed as a “soft” version of data pooling between ˇ$\check{}\mathbf{Z}^{\textsf{ITL}}_{0}$ and ˇ$\check{}\mathbf{Z}^{\textsf{DP}}_{0}$. The following theorem characterizes the performance of ˇ$\check{}\mathbf{Z}^{\lambda}_{0}$, whose proof is deferred to Section B.1.
Theorem 1.
For $s>0$, $\varepsilon\in[0,1/2]$ and $\lambda\geq 0$, define
$\mathcal{M}(s,\varepsilon,\lambda):=\Phi(-\sqrt{2}s)+\Phi\bigg{(}-s-\frac{\lambda}{2s}\bigg{)}\Phi\bigg{(}s-\frac{\lambda}{2s}\bigg{)}+2\varepsilon\Phi\bigg{(}-s+\frac{\lambda}{2s}\bigg{)}.$
The estimate $\check{Z}^{\lambda}_{0,i}$ in (2.2) satisfies
$\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i})\leq\mathcal{M}\left({\mu}/{\sigma},\varepsilon,\lambda\right),\qquad\forall i\in[n].$
When $\varepsilon=0$, labels are perfectly matched and the error is purely caused by noise. In this case, the error bound given by Theorem 1 is $\psi\left(\lambda\right):=\mathcal{M}\left({\mu}/{\sigma},0,\lambda\right)$. When $\varepsilon>0$, the label discrepancy leads to an additional term $\phi\left(\lambda\right):=2\varepsilon\Phi\left(-\mu/\sigma+{\lambda}\left(2\mu/\sigma\right)^{-1}\right)$. Therefore, $\psi\left(\lambda\right)$ and $\phi\left(\lambda\right)$ can be interpreted as stochastic error and systematic error, respectively. Note that $\psi\left(\lambda\right)$ is non-increasing with $\lambda$ and $\phi\left(\lambda\right)$ is non-decreasing with $\lambda$, and the relation $\mathcal{M}\left({\mu}/{\sigma},\varepsilon,\lambda\right)=\psi\left(\lambda\right)+\phi\left(\lambda\right)$ can be viewed as the bias-variance decomposition. In Theorem 3 below, we will show that choosing $\lambda=\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)$ as in (3) optimizes ${\cal M}\left(\mu/\sigma,\varepsilon,\lambda\right)$ up to a constant factor, and achieve the bias-variance trade-off. Formally, we have the following theorem regarding ˇ$\check{}\mathbf{Z}_{0}^{\lambda}$, whose proof can be found in Appendix B.2.
Theorem 2.
Define
$\alpha:=\frac{\log\left(1/\varepsilon\right)}{4\textsf{SNR}},$
then there exists some universal constant $C_{0}>0$ such that
$\inf_{\lambda\geq 0}\mathbb{E}\ell\left(\check{}\mathbf{Z}^{\lambda}_{0},\mathbf{Z}^{*}_{0}\right)\leq C_{0}\exp\left(-\textsf{SNR}\cdot\min\left\{\left(1+\alpha\right)^{2},2\right\}\right).$
By Theorem 2, we can see that $\alpha$ characterizes the informativeness of source data. Denote by $\lambda^{*}$ the quantity such that $\mathbb{E}\ell\left(\check{}\mathbf{Z}^{\lambda^{*}}_{0},\mathbf{Z}^{*}_{0}\right)\asymp\inf_{\lambda\geq 0}\mathbb{E}\ell\left(\check{}\mathbf{Z}^{\lambda}_{0},\mathbf{Z}^{*}_{0}\right)$. In view of Proposition 1, we have
$\mathbb{E}\ell\left(\check{}\mathbf{Z}^{\textsf{ITL}}_{0},\mathbf{Z}^{*}_{0}\right)\lesssim\exp\left(-\textsf{SNR}\right)\qquad\text{and}\qquad\mathbb{E}\ell\left(\check{}\mathbf{Z}^{\textsf{DP}}_{0},\mathbf{Z}^{*}_{0}\right)\lesssim\exp\left(-\textsf{SNR}\cdot\min\left\{4\alpha,2\right\}\right).$
We can thereby compare the performance of ˇ$\check{}\mathbf{Z}^{\lambda^{*}}$ with the ITL estimator ˇ$\check{}\mathbf{Z}^{\textsf{ITL}}$ and the DP estimator ˇ$\check{}\mathbf{Z}^{\textsf{DP}}$, summarized in Table 1. For simplicity, we consider the regime $\textsf{SNR}=o\left(\log n\right)$.
  • ˇ$\check{}\mathbf{Z}^{\lambda^{*}}_{0}$ beats ˇ$\check{}\mathbf{Z}^{\textsf{ITL}}_{0}$ for $\alpha\in(0,\infty]$; ˇ$\check{}\mathbf{Z}^{\lambda^{*}}_{0}$ beats ˇ$\check{}\mathbf{Z}^{\textsf{DP}}_{0}$ for $\alpha\in(0,1/2)$, and performs as well as ˇ$\check{}\mathbf{Z}^{\textsf{DP}}_{0}$ for $\alpha\in[1/2,\infty)$.
  • When $\alpha={\sqrt{2}-1}$, ˇ$\check{}\mathbf{Z}^{\lambda^{*}}_{0}$ can achieve a rate of $\exp\left(-2\cdot\textsf{SNR}\left(1+o(1)\right)\right)$ (indicated by Proposition 1), which is an ideal rate when there is no mismatch ($\varepsilon=0$), while ˇ$\check{}\mathbf{Z}^{\textsf{DP}}_{0}$ can only achieve a rate of
    $\exp\left(-4\left(\sqrt{2}-1\right)\cdot\textsf{SNR}\left(1+o(1)\right)\right)\approx\exp\left(-1.66\cdot\textsf{SNR}\left(1+o(1)\right)\right).$
Range of $\varepsilon$ $<e^{-2\textsf{SNR}}$ $e^{-2\textsf{SNR}}\sim e^{-4\left(\sqrt{2}-1\right)\textsf{SNR}}$ $>e^{-4\left(\sqrt{2}-1\right)\textsf{SNR}}$
Range of $\alpha$ $[1/2,\infty)$ $[\sqrt{2}-1,1/2)$ $(0,\sqrt{2}-1)$
TC optimal as if $\varepsilon=0$ optimal as if $\varepsilon=0$ optimal
DP optimal as if $\varepsilon=0$ sub-optimal sub-optimal
ITL sub-optimal sub-optimal sub-optimal
Table 1: Comparison of TC, DP and ITL under the regime $\textsf{SNR}=o\left(\log n\right)$.
Remark 1 (Comparison with testing).
In [19], the authors proposed a testing procedure for the following hypothesis testing problem:
$H_{0}:\ell\left(\mathbf{Z}_{0}^{*},\mathbf{Z}_{1}^{*}\right)=0\quad\text{vs.}\quad H_{1}:\ell\left(\mathbf{Z}_{0}^{*},\mathbf{Z}_{1}^{*}\right)>\varepsilon$ (5)
for some $\varepsilon\geq 0$, under the regime $\varepsilon=n^{-\beta}$ and $\mu/\sigma=\sqrt{r\log n}$ with some constants $\beta,r>0$. In particular (adapted to our notation), they derived a sharp boundary for (5) in the sense that the likelihood ratio test and a Higher-Criticism-type test [16] is consistent when $\alpha<\alpha^{*}\left(r\right)$, and no test is consistent when $\alpha>\alpha^{*}(r)$, where
$\alpha^{*}(r):=\begin{cases}\frac{1}{4}\left(3+\frac{1}{r}\right),&0<r\leq 1/5\\ \frac{1}{2r}\sqrt{1-\left(1-2r\right)^{2}_{+}},&r>1/5.\end{cases}$
  • When $20<\alpha<\alpha^{*}(r)\wedge 1/$ 2 (Region ${\cal R}_{1,\sf det}$ in Figure 1), $\varepsilon$ is sufficiently large so that we can tell the the difference between $\mathbf{Z}_{0}^{*}$ and $\mathbf{Z}_{1}^{*}$. A natural strategy is to use independent task learning estimator solely based on $\mathbf{X}_{0}$, leading to an error rate of $\exp\left(-\textsf{SNR}\left(1+o(1)\right)\right)$. On the other hand, applying data pooling leads to an error rate of $\exp\left(-4\alpha\cdot\textsf{SNR}\left(1+o(1)\right)\right)$. In contrast, the transfer learning with optimal $\lambda$ leads to a vanishingly smaller error rate $\exp\left(-(1+\alpha)^{2}\cdot\textsf{SNR}\left(1+o(1)\right)\right)$.
  • When $1/2<\alpha<\alpha^{*}(r)\wedge 1$ (Region ${\cal R}_{2,\sf det}$ in Figure 1), $\varepsilon$ remains large enough to be detectable. Here, independent task learning estimator is still sub-optimal. The transfer learning with optimal $\lambda$ achieves the ideal error rate $\exp\left(-2\cdot\textsf{SNR}\left(1+o(1)\right)\right)$, which is also attainable by data pooling. This regime reveals a subtle difference between clustering and testing: even when the discrepancy is detectable, the optimal estimator can be constructed by ignoring it.
  • When $\alpha^{*}(r)<\alpha<1$ (See Region ${\cal R}_{\sf undet}$ in Figure 1), $\varepsilon$ is so small that no test can reliably tell the difference between $\mathbf{Z}_{0}^{*}$ and $\mathbf{Z}_{1}^{*}$. By definition of $\alpha$ and $\alpha^{*}\left(r\right)$, we have $\beta>1$ in this regime. Both the transfer learning with optimal $\lambda$ and data pooling result in the error rates of $\exp\left(-\min\left\{2,{2}/{r}\right\}\cdot\textsf{SNR}\left(1+o(1)\right)\right)$.
Figure 1: Comparison with the hypothesis test’s detection boundary. Below the red curve $\alpha=\alpha^{*}(r)$, i.e., Region ${\cal R}_{1,\sf det}\bigcup{\cal R}_{2,\sf det}$, consistent test exists for (5); above the red curve, i.e., Region ${\cal R}_{\sf undet}$, there is no consistent test for (5). Region ${\cal R}_{\sf undet}$: transfer clustering performs as well as data pooling when no consistent test can distinguish $\mathbf{Z}_{0}^{*}$ and $\mathbf{Z}_{1}^{*}$. Region ${\cal R}_{1,\sf det}$: transfer clustering beats data pooling while we can have a consistent test to distinguish $\mathbf{Z}_{0}^{*}$ and $\mathbf{Z}_{1}^{*}$. Region ${\cal R}_{2,\sf det}$: transfer clustering performs as well as data pooling while we can have a consistent test to distinguish $\mathbf{Z}_{0}^{*}$ and $\mathbf{Z}_{1}^{*}$.
We certify the optimality of our proposed transfer learning procedure (2.2) by establishing a matching lower bound for estimating $\mathbf{Z}_{0}^{*}$. The proof is deferred to Appendix B.3.
Theorem 3 (Lower bound).
Assume that $\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i}|Z^{*}_{0,i})=\varepsilon$. For any estimate ^$\widehat{}\mathbf{Z}_{0}$ of $\mathbf{Z}_{0}^{*}$, we have
$\mathbb{P}(\widehat{Z}_{0,i}\neq Z_{0,i}^{*})\geq\frac{1}{4}\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right),\qquad\forall i\in[n].$
where the function $\mathcal{M}$ is defined in Theorem 1. The right-hand side also lower bounds $\mathbb{E}\ell(\widehat{}\mathbf{Z}_{0},\mathbf{Z}_{0}^{*})$.
2.3 Adaptivity: Goldenshluger-Lepski method
In view of Theorem 1 and Theorem 3, the optimal choice of $\lambda$ for our TC estimator should be $\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)$. The key challenge we are going to tackle is the unknown $\varepsilon$. For clarity of presentation, we assume $\left(\mu,\sigma\right)$ is given by an oracle. In principle, estimating $\left(\mu,\sigma\right)$ is much easier than estimating $\varepsilon$.
A direct method is to first estimate $\varepsilon$ by $\widehat{\varepsilon}:=\ell\left(\check{}\mathbf{Z}^{\textsf{ITL}}_{0},\check{}\mathbf{Z}^{\textsf{ITL}}_{1}\right)$ and then use $\log\left(\widehat{\varepsilon}^{-1}\left(1-\widehat{\varepsilon}\right)\right)$ as a plug-in estimator of $\log\left(\widehat{\varepsilon}^{-1}\left(1-\widehat{\varepsilon}\right)\right)$. However, the following lemma shows that the direct plug-in method may not work for $\varepsilon$ when $\varepsilon$ is very small [^2] A possible approach for estimating small $\varepsilon$ for GMM is given by [7], which, however, heavily relies on the Gaussianity assumption in the model. Our method can be easily adapted to other distributions, see Algorithm 2 and Algorithm 3. :
Lemma 1.
Assume that $\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i}|Z^{*}_{0,i})=\varepsilon=n^{-\beta}$ for $\beta\in(0,1)$. If
$\beta>\min\left\{\frac{\textsf{SNR}}{\log n},~{}\frac{1}{2}\left(\frac{\textsf{SNR}}{\log n}+1\right)\right\},$
then we have
$\mathbb{E}\left(\frac{\widehat{\varepsilon}}{\varepsilon}-1\right)^{2}\geq c_{0},$
for some universal constant $c_{0}>0$.
Consider the case when $\textsf{SNR}=o\left(\log n\right)$, then Lemma 1 tells us $\widehat{\varepsilon}$ fails to consistently estimate $\varepsilon$ if $\varepsilon=o\left(n^{-1/2}\right)$. Consequently, $\log\left(\widehat{\varepsilon}^{-1}\left(1-\widehat{\varepsilon}\right)\right)$ is a not consistent estimator of $\log\left(\varepsilon\left(1-\varepsilon\right)\right)$. To bridge this gap, we develop a novel adaptive method that bypasses the estimation of $\varepsilon$ to choose $\lambda$, inspired by the Goldenshluger-Lepski method [24]. The high-level idea is outlined below.
Consider a grid set $\Lambda=\left\{0=\lambda_{1}<\lambda_{2}<\cdots<\lambda_{M}=\log n\right\}$. Recall the bias-variance decomposition $\mathcal{M}\left({\mu}/{\sigma},\varepsilon,\lambda\right)=\psi\left(\lambda\right)+\phi\left(\lambda\right)$, the key to our adaptive procedure is to construct estimates of $\psi$ and $\phi$ on $\Lambda$. To estimate $\psi$, it boils down to characterizing the stochastic error purely incurred by the noise, i.e., ${\cal M}\left(\mu/\sigma,0,\lambda\right)$. Therefore, we need to construct samples from the distribution where $\varepsilon=0$. Note that we can write $\mathbf{X}_{0}=\mu\mathbf{Z}^{*}_{0}+\mathbf{E}_{0}$ and $\mathbf{X}_{1}=\mu\mathbf{Z}^{*}_{1}+\mathbf{E}_{1}$, where $\mathbf{E}_{0},\mathbf{E}_{1}$ have i.i.d. $N(0,\sigma^{2})$ entries. Suppose we have access to an imaginary sample ¯$\bar{}\mathbf{X}_{0}=\mu\mathbf{Z}^{*}_{0}+\mathbf{E}_{1}$, and let (¯$(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\lambda}_{1})=\textsf{TC}\left(\mathbf{X}_{0},\bar{}\mathbf{X}_{0};\lambda,\mu,\sigma\right)$ for $\lambda\in\Lambda\cup\left\{\infty\right\}$. ¯$\bar{}\mathbf{X}_{0}$ is a “hybrid” sample of target and source, in the sense that it has the same label as $\mathbf{X}_{0}$, while the same noise as $\mathbf{X}_{1}$. A benchmark estimator should be ¯$\bar{}\mathbf{Z}^{\infty}_{0}$, as $\mathbf{X}_{0}$ and ¯$\bar{}\mathbf{X}_{0}$ share the same labels and (¯$\ell\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right)$ is the best error rate we can achieve when $\varepsilon=0$. It suffices to characterize (ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}^{\infty}_{0}\right)$ due to the triangle inequality. Moreover, we have
(ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}^{\infty}_{0}\right)\leq\ell\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}^{\infty}_{0}\right)+\ell\left(\check{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}^{\lambda}_{0}\right).$
It can be shown that with high probability, (¯$\ell\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}^{\infty}_{0}\right)\lesssim\psi\left(\lambda\right)$ and (ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}^{\lambda}_{0}\right)\lesssim\phi\left(\lambda\right)$. Since $\left(\mu,\sigma\right)$ is known, we can simulate from the distribution of $\left(\mathbf{X}_{0},\bar{}\mathbf{X}_{0}\right)$, generate bootstrap samples {(¯$\left\{\ell\left(\bar{}\mathbf{Z}^{\lambda,b}_{0},\bar{}\mathbf{Z}^{\infty,b}_{0}\right),b\in[B]\right\}$, and construct an estimator $\widehat{\psi}\left(\lambda\right)$ for $\psi\left(\lambda\right)$ by its sample quantile. This procedure is detailed in Algorithm 1.
To estimate $\phi\left(\lambda\right)$, a natural idea from Goldenshluger-Lepski method [24] is to use the quantity
$\max_{\lambda^{\prime}\in\Lambda,\lambda^{\prime}<\lambda}\left[\ell\left(\check{}\mathbf{Z}_{0}^{\lambda},\check{}\mathbf{Z}_{0}^{\lambda^{\prime}}\right)-\widehat{\psi}\left(\lambda\right)-\widehat{\psi}\left(\lambda^{\prime}\right)\right]_{+},$ (6)
which serves as an estimate of $\phi\left(\lambda\right)$ by eliminating two estimates of stochastic error. The aforementioned G-L method is a powerful tool for adaptive nonparametric estimation. For our purposes, we can further simplify (6) and use the following estimator for $\phi\left(\lambda\right)$:
$\quad\widehat{\phi}\left(\lambda\right):=\max_{\lambda^{\prime}\in\Lambda,\lambda^{\prime}<\lambda}\left[\ell\left(\check{}\mathbf{Z}_{0}^{\lambda},\check{}\mathbf{Z}_{0}^{\lambda^{\prime}}\right)-\widehat{\psi}\left(\lambda^{\prime}\right)\right]_{+}.$
The final choice of $\lambda$ is given by
$\widehat{\lambda}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}\left\{\widehat{\phi}\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right\}.$
This leads to our adaptive estimator
$\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,\mu,\sigma\right)=\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},$
where ATC is a shorthand for Adaptive Transfer Clustering.
Input: Parameters $\mu,\sigma>0$, $\zeta\in(0,1)$.
1 for $b=1,\cdots,B$ do
2       Generate $Z^{*,b}_{0,i}\overset{i.i.d.}{\sim}\text{Unif}\left\{\pm 1\right\}$ for $i\in[n]$.
3      Generate $X_{0,i}^{b}\overset{i.i.d.}{\sim}N\left(Z^{*,b}_{0,i}\mu,\sigma^{2}\right)$ and $X_{1,i}^{b}\overset{i.i.d.}{\sim}N\left(Z^{*,b}_{0,i}\mu,\sigma^{2}\right)$ for $i\in[n]$ independently.
4      For each $\lambda\in\Lambda\cup\left\{\infty\right\}$, compute (¯$\big{(}\bar{}\mathbf{Z}_{0}^{\lambda,b},\bar{}\mathbf{Z}_{1}^{\lambda,b}\big{)}:=\textsf{TC}\left(\mathbf{X}^{b}_{0},\mathbf{X}^{b}_{1};\lambda,\mu,\sigma\right)$.
5 end for
6For each $\lambda\in\Lambda$, compute $\widehat{\psi}\left(\lambda\right):=1.01\widehat{Q}_{\zeta/2}^{\lambda}$, where $\widehat{Q}_{\zeta/2}^{\lambda}$ is $\left(1-\zeta/2\right)$ quantile of {(¯$\left\{\ell\left(\bar{}\mathbf{Z}^{\lambda,b}_{0},\bar{}\mathbf{Z}^{\infty,b}_{0}\right),b\in[B]\right\}$.
Output: $\left\{\widehat{\psi}\left(\lambda\right),\lambda\in\Lambda\right\}$
7
Algorithm 1 Bootstrap-Quantile for $\psi$ in two-component GMM
The following theorem provides an error bound for ATC, whose proof can be found in Appendix B.4.
Theorem 4.
Let $\zeta\in(0,1)$, and assume that $B\geq C_{0}\zeta^{-2}\log\left(M\zeta^{-1}\right)$ for some universal constant $C_{0}>0$, then with probability at least $1-\zeta$, we have
(ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{1}\left[\min_{\lambda\in\Lambda}\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\lambda\right)+\frac{\log\left(M\zeta^{-1}\right)}{n}\right].$
for some universal constant $C_{1}>0$ determined by $C_{0}$.
Due to the smoothness of $\Phi$, it follows that that ATC can achieve the optimal clustering error rate $\mathcal{M}\left({\mu}/{\sigma},\varepsilon,\log\left({\varepsilon}^{-1}\left(1-\varepsilon\right)\right)\right)$, up to a negligible additive term ${\log(M\zeta^{-1})}/n$, provided that the grid in $\Lambda$ is sufficiently dense. Formally, we have the following result, whose proof can be found in Appendix C.1.
Corollary 1.
Suppose the conditions of Theorem 4 hold. Let $M=\lceil\log^{2}n\rceil$ and $\lambda_{j}=jM^{-1}\log n$ for $j=0,\cdots,M$, then there exists some universal constant $C_{2}>0$ determined by $C_{1}$ such that with probability at least $1-\zeta$,
(ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{2}\left[\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right)+\frac{\log\left(M\zeta^{-1}\right)}{n}\right].$
Moreover, if $\textsf{SNR}<\left(\frac{1}{2}-c\right)\log n$ for some constant $c\in(0,1)$, then there exists some universal constant $C_{3}>0$ determined by $C_{2}$ such that with probability at least $1-n^{-10}$,
(ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{3}\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right).$ (7)
Corollary 1 delivers that ATC estimator achieves the error rate that matches the lower bound in Theorem 3.
3 Methodology for transfer clustering in general mixture models
In this section, we present our general framework for transfer learning in mixture models.
3.1 Problem setup
In the two-component symmetric GMM, our data can be written as $\mathbf{X}_{0}=\mu\mathbf{Z}_{0}^{*}+\sigma\mathbf{E}_{0}$ and $\mathbf{X}_{1}=\mu\mathbf{Z}_{1}^{*}+\sigma\mathbf{E}_{1}$, where $\mathbf{E}_{0}$ and $\mathbf{E}_{1}$ are independent $N\left(0,\mathbf{I}_{n}\right)$. We now extend it to a more general data generating mechanism. Consider two generative models ${\cal G}_{0}$ and ${\cal G}_{1}$ that generate the target data $X$ and source data $Y$ respectively:
$\mathbf{X}_{0}={\cal G}_{0}\left(\mathbf{Z}_{0}^{*},\mathbf{\Theta}_{0}^{*},\mathbf{E}_{0}\right),\quad\mathbf{X}_{1}={\cal G}_{1}\left(\mathbf{Z}_{1}^{*},\mathbf{\Theta}_{1}^{*},\mathbf{E}_{1}\right),$
where $\mathbf{Z}_{0}^{*},\mathbf{Z}_{1}^{*}\in[K]^{n}$ are latent labels of our interest, $\mathbf{\Theta}_{0}^{*}$ and $\mathbf{\Theta}_{1}^{*}$ are nuisance parameters, and $\mathbf{E}_{0}$ and $\mathbf{E}_{1}$ are random seeds. Our goal is to estimate $\mathbf{Z}_{0}^{*}$, whose performance is measured by
${\mathfrak{D}}(\mathbf{Z},\mathbf{Z}^{\prime}):=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left(Z_{i}\neq Z^{\prime}_{i}\right),\qquad\forall\mathbf{Z},\mathbf{Z}^{\prime}\in[K]^{n}.$
Denote by $L_{m}\left(\mathbf{Z},\mathbf{X};\mathbf{\Theta}\right)$ the posterior density of either target data ($m=0$) or source data ($m=1$). We list several illustrative examples on the generating mechanism of target and source data.
Example 3.1 (Gaussian mixture model, warm-up case in Section 2).
A primary example of theoretical interest is the two-component symmetric Gaussian mixture model discussed in Section 2. For $m\in\left\{0,1\right\}$, we have $\mathbf{\Theta}_{m}^{*}=\left(\mu,\sigma\right)$, and ${\cal G}_{m}\left(\mathbf{Z},\left(\mu,\sigma\right),\mathbf{E}\right)=\mu\mathbf{Z}+\sigma\mathbf{E}$. We thus have $L_{0}=L_{1}=L_{\textsf{GMM}_{2}}$ with
$\log L_{\textsf{GMM}_{2}}\left(\mathbf{Z},\mathbf{X};\left(\mu,\sigma\right)\right)=\sum_{i=1}^{n}\left[\frac{(X_{0,i}-\mu)^{2}}{2\sigma^{2}}+\log\left(1/2\right)\right].$
Example 3.2 (Gaussian mixture model, general case).
We can extend the warm-up case to the general scenario as follows. In this case for $m\in\left\{0,1\right\}$, we have $\mathbf{\Theta}_{m}^{*}=\left(\left\{\pi_{m,k}\right\}_{k=1}^{K},\left\{{\bm{\mu}}_{m,k}\right\}_{k=1}^{K},\left\{\mathbf{\Sigma}_{m,k}\right\}_{k=1}^{K}\right)$, $\mathbf{E}_{m}\in\mathbb{R}^{n\times d}$ has i.i.d. $N(0,1)$ entries, and ${\cal G}_{m}\left(\mathbf{Z},\left(\left\{\pi_{k}\right\}_{k=1}^{K},\left\{{\bm{\mu}}_{k}\right\}_{k=1}^{K},\left\{\mathbf{\Sigma}_{k}\right\}_{k=1}^{K}\right),\mathbf{E}\right)\in\mathbb{R}^{n\times d}$ with its $i$-th row being ${\bm{\mu}}_{Z_{i}}+\mathbf{\Sigma}^{1/2}_{Z_{i}}\mathbf{E}_{i,\cdot}^{\top}$. We thus have $L_{0}=L_{1}=L_{\textsf{GMM}_{K}}$ with
$\log L_{\textsf{GMM}_{K}}\left(\mathbf{Z},\mathbf{X};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\left\{{\bm{\mu}}_{k}\right\}_{k=1}^{K},\left\{\mathbf{\Sigma}_{k}\right\}_{k=1}^{K}\right)\right):=\sum_{i=1}^{n}\left[\log{\cal N}\left(\mathbf{X}_{i}\mid{\bm{\mu}}_{Z_{i}},\mathbf{\Sigma}_{Z_{i}}\right)+\log\pi_{Z_{i}}\right],$
where ${\cal N}\left(\cdot\mid{\bm{\mu}},\mathbf{\Sigma}\right)$ is the multivariate normal density with mean ${\bm{\mu}}$ and covariance $\mathbf{\Sigma}$.
Example 3.3 (Latent class model).
Another example is the latent class model [LCM; 26, 38], which is a popular mixture model for multivariate categorical/binary data, widely used in social sciences [66], genetics and genomics [36], and health sciences [67]. Let $m\in\left\{0,1\right\}$. The data $\mathbf{X}_{m}\in\left\{0,1\right\}^{n\times d}$ consists of $d$-dimensional binary response vectors of $n$ subjects. $\left\{\mathbf{X}_{m,i}\in\left\{0,1\right\}^{d}\right\}_{i=1}^{n}$ are assumed to be conditionally independent given the latent label vector $\mathbf{Z}_{m}^{*}$. Let $\mathbf{P}_{m}\in[0,1]^{d\times K}$ be the item parameter matrix, we have $\mathbb{P}\left(X_{m,ij}=1\mid Z^{*}_{m,i}\right)=(\mathbf{P}_{m})_{jZ^{*}_{m,i}}$ for $i\in[n]$ and $j\in[d]$. In addition, we assume entries of $\mathbf{Z}_{m}^{*}$ are i.i.d. categorical random variables with parameters $\left\{\pi_{m,k}\right\}_{k=1}^{K}$ for $m\in\left\{0,1\right\}$. In this case, we have $\mathbf{\Theta}_{m}^{*}=\left(\left\{\pi_{m,k}\right\}_{k=1}^{K},\mathbf{P}_{m}\right)$, $\mathbf{E}_{m}\in[0,1]^{n\times d}$ has i.i.d. Unif$[0,1]$ entries, and we can write ${\cal G}_{m}\left(\mathbf{Z},\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{P}\right),\mathbf{E}\right)\in\left\{0,1\right\}^{n\times d}$ with its $(i,j)$-th entry being $\mathbb{I}\left(E_{ij}\leq P_{jZ_{i}}\right)$. We thus have $L_{0}=L_{1}=L_{\textsf{LCM}}$ with
$\log L_{\textsf{LCM}}$ $\left(\mathbf{Z},\mathbf{X};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{P}\right)\right)$
$:=\sum_{i=1}^{n}\sum_{j=1}^{d}\left[X_{ij}\log P_{jZ_{i}}+\left(1-X_{ij}\right)\log\left(1-P_{jZ_{i}}\right)\right]+\sum_{i=1}^{n}\log\pi_{Z_{i}}.$
Example 3.4 (Contextual Stochastic Block Model).
Our general framework also fits the contextual SBM [43, 14]. We first introduce SBM, which is probably the most well-studied model in thes network literature [28, 1]. Let $\mathbf{X}_{0}\in\left\{0,1\right\}^{n\times n}$ be the adjacency matrix for a graph of $n$ nodes, $\mathbf{Z}_{0}^{*}\in[K]^{n}$ be the community label vector and $\mathbf{B}\in[0,1]^{K\times K}$ be the link matrix. For a symmetric SBM, we assume $\mathbb{P}\left(X_{0,ij}=1\mid\mathbf{B},\mathbf{Z}^{*}_{0}\right)=B_{Z^{*}_{i}Z^{*}_{j}}$ and $X_{ij}=X_{ji}$ and $X_{ii}=0$ for $i\neq j\in[n]$. Moreover, $\left\{\mathbf{X}_{0,ij}\right\}_{i<j}$ are conditionally independent given $\mathbf{Z}^{*}_{0}$ and $\mathbf{B}$, and entries of $\mathbf{Z}_{0}^{*}$ are i.i.d. categorical random variables with parameters $\left\{\pi_{0,k}\right\}_{k=1}^{K}$. We have $\mathbf{\Theta}_{0}^{*}=\left(\left\{\pi_{0,k}\right\}_{k=1}^{K},\mathbf{B}\right)$, and ${\cal G}_{0}$ can be defined similarly to LCMs. The posterior $L_{\textsf{SBM}}$ is defined as
$\log L_{\textsf{SBM}}$ $\left(\mathbf{Z},\mathbf{X};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{B}\right)\right)$
$:=\sum_{1\leq i<j\leq n}\left[X_{ij}\log B_{Z_{i}Z_{j}}+\left(1-X_{ij}\right)\log\left(1-B_{Z_{i}Z_{j}}\right)\right]+\sum_{i=1}^{n}\log\pi_{Z_{i}}.$
In the contextual setting, in addition we have $\left\{\mathbf{X}_{1,i}\in\mathbb{R}^{d}\right\}_{i=1}^{n}$ representing the features of nodes, which follow the Gaussian mixture model with parameter $\mathbf{\Theta}_{1}^{*}=\left(\left\{\pi_{1,k}\right\}_{k=1}^{K},\left\{{\bm{\mu}}_{k}\right\}_{k=1}^{K},\left\{\mathbf{\Sigma}_{k}\right\}_{k=1}^{K}\right)$. We thus have $L_{0}=L_{\sf SBM}$, $L_{1}=L_{{\sf GMM}_{K}}$. It is worth noting that we can exchange the roles of target and source data to have $L_{0}=L_{{\sf GMM}_{K}}$ and $L_{1}=L_{\sf SBM}$, i.e., we treat the feature $\mathbf{X}_{1}$ as target and use network $\mathbf{X}_{0}$ as source to help us estimate $\mathbf{Z}_{1}^{*}$, which is the latent label vector of $\mathbf{X}_{1}$.
3.2 General transfer clustering methodology
We will develop an adaptive transfer clustering method for the general framework. For any nuisance parameter (^$(\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1})$, our transfer learning estimators for the clustering labels are defined as follows:
(^$\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\widehat{}\mathbf{Z}_{1}^{\lambda}\right)=\operatorname*{arg\,min}_{\mathbf{Z}_{0},\mathbf{Z}_{1}\in[K]^{n}}\left\{-\log L_{0}\left(\mathbf{Z}_{0},\mathbf{X}_{0};\widehat{}\mathbf{\Theta}_{0}\right)-\log L_{1}\left(\mathbf{Z}_{1},\mathbf{X}_{1};\widehat{}\mathbf{\Theta}_{1}\right)+\lambda n{\mathfrak{D}}\left(\mathbf{Z}_{0},\mathbf{Z}_{1}\right)\right\},$ (8)
where $\lambda>0$ is a tuning parameter. The procedure (8), with slight abuse of notation, is denoted by $\textsf{TC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\lambda,\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1}\right)$. Here, instead of assuming known parameters in warm-up example, we need to have an estimator for nuisance parameters $(\mathbf{\Theta}_{0},\mathbf{\Theta}_{1})$ (See 2). In LCMs and related latent variable models, this corresponds to the common item parameter calibration setting [34, 57], where the parameters have been estimated and calibrated accurately with previous samples.
Given a grid set $\Lambda\subset[0,\infty)$ and $\widehat{\psi}$ on $\Lambda$, we generalize the Goldenshluger-Lepski procedure in Section 2 to Algorithm 2, in order to select the optimal $\lambda$. Here $\widehat{\psi}$ is estimated by using a bootstrap procedure detailed in Algorithm 3, which can be regarded as a generalization to Algorithm 1. The final estimator is denoted by
$\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1}\right)=\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}}.$
A natural and commonly asked question is why not directly estimate a joint model with both the source and target data, where $[K]\times[K]$ is the label space for $(Z_{0,i},Z_{1,i})$. While this might seem like a straightforward approach in the current literature [22, 23], there are several reasons why it is less favorable than our proposed method. Firstly, when using a joint model with $[K]\times[K]$ labels, the number of clusters increases significantly to $K^{2}$, which can be computationally inefficient. In the interesting regime where $\varepsilon=o(1)$, the sizes of these clusters become extremely unbalanced, making it more difficult to achieve high statistical accuracy. Joint modeling is necessary in [22, 23], as their focus is on testing the independence of clusters between target and source data. In contrast, our method aims to improve clustering performance of the target data by estimating only $K$ clusters in the target domain, and using a discrepancy parameter $\varepsilon$ to gauge the discrepancy between the source and target domains. This avoids overfitting to the noise present in smaller clusters. Secondly, a joint model can lead to difficulties in fitting when combining different types of models for the source and target datasets, e.g., contextual SBM. Our approach circumvents this complexity by separately fitting the models for the source and target data, then aggregating them adaptively. This makes our approach easily adaptable to different data structures and combinable with existing estimators.
Remark 2.
Our methodology framework can be readily adapted to other latent variable models beyond clustering (e.g., mixed membership models, factor models). For latent variables $\mathbf{Z}_{0}^{*},\mathbf{Z}_{1}^{*}\in{\cal Z}^{n}$ of our interest, where ${\cal Z}$ is some metric space, we only need to introduce an appropriate metric in ${\cal Z}^{n}$ in place of ${\mathfrak{D}}$, and modify the penalty term in (8) accordingly.
Input: Data $(\mathbf{X}_{0},\mathbf{X}_{1})$, estimated parameter (^$(\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1})$, grid set $\Lambda$, function $\widehat{\psi}$ on $\Lambda$
1 For each $\lambda\in\Lambda$, compute ^$\widehat{}\mathbf{Z}_{0}^{\lambda}$ according to (8).
2For each $\lambda\in\Lambda$, compute $\widehat{\phi}\left(\lambda\right)=\max_{\lambda^{\prime}\in\Lambda,\lambda^{\prime}<\lambda}\left[{\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\widehat{}\mathbf{Z}_{0}^{\lambda^{\prime}}\right)-\widehat{\psi}\left(\lambda^{\prime}\right)\right]_{+}$.
3Let $\widehat{\lambda}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}\left\{\widehat{\phi}\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right\}$.
Output: $\widehat{\lambda}$
Algorithm 2 Goldenshluger-Lepski method
Input: Estimated parameter (^$(\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1})$, tuning parameter $\zeta\in(0,1)$
1 for $b=1,\cdots,B$ do
2       Generate $Z^{*,b}_{0,i}\overset{i.i.d.}{\sim}\widehat{\Pi}_{0}$ for $i\in[n]$, where $\widehat{\Pi}_{0}$ is the estimated parameter for marginal distribution of $\mathbf{Z}_{0}$ included in ^$\widehat{}\mathbf{\Theta}_{0}$.
3      Generate $\mathbf{X}_{0}^{b}={\cal G}_{0}\left(\mathbf{Z}_{0}^{*,b},\widehat{}\mathbf{\Theta}_{0},\mathbf{E}_{0}^{b}\right)$ and $\mathbf{X}_{1}^{b}={\cal G}_{1}\left(\mathbf{Z}_{0}^{*,b},\widehat{}\mathbf{\Theta}_{1},\mathbf{E}_{1}^{b}\right)$.
4      For each $\lambda\in\Lambda\cup\left\{\infty\right\}$, compute (¯$\left(\bar{}\mathbf{Z}^{\lambda,b}_{0},\bar{}\mathbf{Z}^{\lambda,b}_{1}\right)=\textsf{TC}\left(\mathbf{X}^{b}_{0},\mathbf{X}^{b}_{1};\lambda,\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1}\right)$.
5 end for
6For each $\lambda\in\Lambda$, compute $\widehat{\psi}\left(\lambda\right)=1.01\widehat{Q}_{\zeta/2}^{\lambda}$, where $\widehat{Q}_{\zeta/2}^{\lambda}$ is $\left(1-\zeta/2\right)$ quantile of $\left\{{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda,b}_{0},\bar{}\mathbf{Z}^{\infty,b}_{0}\right),b=[B]\right\}$.
Output: $\left\{\widehat{\psi}\left(\lambda\right),\lambda\in\Lambda\right\}$
Algorithm 3 Bootstrap-Quantile for $\psi$ in general mixture models
4 General theory
In this section, we conduct a theoretical analysis of the adaptive estimator $\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1}\right)$ based on Algorithm 2 and Algorithm 3. We start with general results and then derive sharp adaptivity guarantees under the two-component Gaussian mixture model in $\mathbb{R}^{d}$.
4.1 Assumptions and general theories
We impose the following assumptions on $\left(\mathbf{Z}^{*}_{0},\mathbf{Z}^{*}_{1}\right)$:
  • ${Z^{*}_{0,i}}\overset{i.i.d.}{\sim}\Pi_{0}$ and ${Z^{*}_{1,i}}\overset{i.i.d.}{\sim}\Pi_{1}$ for all $i\in[n]$, where $\Pi_{0}$ and $\Pi_{1}$ are probability distributions over $[K]$;
  • $\mathbb{P}\left(Z^{*}_{0,i}\neq Z^{*}_{1,i}\right)\leq\varepsilon$ for $i\in[n]$.
For technical reasons, we assume the logarithmic posterior densities have separable forms:
Assumption 1 (Separability).
For $m\in\left\{0,1\right\}$, there exists $\left\{L_{m,i}\right\}_{i=1}^{n}$ such that $\log L_{m}\left(\mathbf{Z},\mathbf{X};\mathbf{\Theta}\right)=\sum_{i=1}^{n}\log L_{m,i}\left(Z_{i},X_{i};\mathbf{\Theta}\right)$.
Remark 3.
Note that Assumption 1 fails for the posterior of the SBM. In particular, we have $\log L_{\textsf{SBM}}\left(\mathbf{Z},\mathbf{X};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{B}\right)\right)=\sum_{i=1}^{n}\log L_{\textsf{SBM},i}\left(Z_{i},X_{i};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{B},\mathbf{Z}_{-i},\mathbf{X}_{-i}\right)\right)$, where $\mathbf{Z}_{-i}\in[K]^{N-1}$ is obtained by removing the $i$-th entry of $\mathbf{Z}$, $\mathbf{X}_{-i}\in\mathbb{R}^{(n-1)\times(n-1)}$ is obtained by removing the $i$-th row and column of $\mathbf{X}$, and
$\log L_{\textsf{SBM},i}$ $\left(Z_{i},X_{i};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{B},\mathbf{Z}_{-i},\mathbf{X}_{-i}\right)\right)$
$:=\sum_{j>i}\left[X_{ij}\log B_{Z_{i}Z_{j}}+\left(1-X_{ij}\right)\log\left(1-B_{Z_{i}Z_{j}}\right)\right]+\log\pi_{Z_{i}}.$
However, the optimization (8) can still be decoupled across different $i\in[n]$ by using $L_{\textsf{SBM},i}\left(Z_{i},X_{i};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{B},\mathbf{Z}_{-i},\mathbf{X}_{-i}\right)\right)$, in which $\mathbf{Z}_{-i}$ and $\mathbf{X}_{-i}$ are viewed as part of parameters. This approach is commonly used in the network analysis literature [2, 3, 20].
Consider an imaginary source data ¯$\bar{}\mathbf{X}_{0}={\cal G}_{1}\left(\mathbf{Z}_{0}^{*},\mathbf{\Theta}^{*}_{1},\mathbf{E}_{1}\right)$, and define
(¯$\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{1}^{\lambda}\right):=\textsf{TC}\left(\mathbf{X}_{0},\bar{}\mathbf{X}_{0};\lambda,\mathbf{\Theta}^{*}_{0},\mathbf{\Theta}^{*}_{1}\right).$
We can view ¯$\bar{}\mathbf{X}_{0}$ as the source data generated by the “clean” model ${\cal G}_{1}\left(\mathbf{Z}_{0}^{*},\mathbf{\Theta}^{*}_{1},\mathbf{E}_{1}\right)$, as it shares the same latent labels with target data $\mathbf{X}_{0}$. Observe that
  • ¯$\bar{}\mathbf{Z}_{0}^{0}$ reduces to the independent task learning estimator using only $\mathbf{X}_{0}$.
  • ¯$\bar{}\mathbf{Z}_{0}^{\infty}$ reduces to the data-pooling estimator, which can also be regarded as the MAP under the “null”, i.e., $\varepsilon=0$.
It is thereby natural to compare ${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\mathbf{Z}_{0}^{*}\right)$ to ${\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\infty},\mathbf{Z}_{0}^{*}\right)$, for which we get
$\left|{\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\mathbf{Z}_{0}^{*}\right)-{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\infty},\mathbf{Z}_{0}^{*}\right)\right|\leq{\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq\underbrace{{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)}_{\textsf{``variance"}}+\underbrace{{\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)}_{\textsf{``bias"}}$
Note that the variance term ${\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ only involves a “clean” model (i.e., with perfectly matched cluster labels) and can be estimated by parametric bootstrap if $\mathbf{\Theta}^{*}_{0}$ and $\mathbf{\Theta}^{*}_{1}$ are known. Define (ˇ$\left(\check{}\mathbf{Z}_{0}^{\lambda},\check{}\mathbf{Z}_{1}^{\lambda}\right):=\textsf{TC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\lambda,\mathbf{\Theta}^{*}_{0},\mathbf{\Theta}^{*}_{1}\right)$. It can be readily seen that the bias term ${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)$ can be decomposed as
[…]
Ideally, for Algorithm 2, we want an oracle inequality of the form
${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)\lesssim\min_{\lambda\in\Lambda}\{\phi(\lambda)+\psi(\lambda)\}+\text{overhead}.$
We present the following theorem of the same flavor. See Section D.2 for the proof.
Lemma 2 (Oracle inequality for Goldenshluger-Lepski (GL) method).
Consider any set $\Lambda$ and $\widehat{\psi}$ on $\Lambda$, let $\widehat{\lambda}$ be obtained from Algorithm 2. Suppose there is a deterministic function $\phi$ on $\Lambda$ such that
  1. 1.
    $\mathbb{P}\left({\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)\leq\phi\left(\lambda\right),\forall\lambda\in\Lambda\right)\geq 1-\zeta$,
  2. 2.
    $\phi$ is non-decreasing.
Then the following inequality holds with probability at least $1-\zeta$:
${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq 4\min_{\lambda\in\Lambda}\left\{\phi\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right\}+3\xi_{\widehat{\psi}},$
where
$\xi_{\widehat{\psi}}:=\max_{\lambda\in\Lambda}\left[{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)-\widehat{\psi}\left(\lambda\right)\right]_{+}.$
If ${\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq\widehat{\psi}\left(\lambda\right)$ for all $\lambda\in\Lambda$, then $\xi_{\widehat{\psi}}=0$ and
${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq 4\min_{\lambda\in\Lambda}\left\{\phi\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right\}.$ (9)
In general, we would hope that ${\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ is dominated by $\widehat{\psi}\left(\lambda\right)$ plus a small quantity, so that $\xi_{\widehat{\psi}}$ is small. This requires $\widehat{\psi}$ to be sufficiently large. Meanwhile, we do not want it to be too large because that would make the upper bound in (9) become vacuous. Fortunately, choosing $\widehat{\psi}$ by quantiles of bootstrap samples does a good job in reconciling the two competing forces.
Define ~$\widetilde{}\mathbf{X}_{0}={\cal G}_{0}\left(\mathbf{Z}_{0}^{*},\widehat{}\mathbf{\Theta}_{0},\mathbf{E}_{0}\right)$, ~$\widetilde{}\mathbf{X}_{1}={\cal G}_{1}\left(\mathbf{Z}_{0}^{*},\widehat{}\mathbf{\Theta}_{1},\mathbf{E}_{1}\right)$ and let
(~$\left(\widetilde{}\mathbf{Z}^{\lambda}_{0},\widetilde{}\mathbf{Z}^{\lambda}_{1}\right):=\textsf{TC}\left(\widetilde{}\mathbf{X}_{0},\widetilde{}\mathbf{X}_{1};\lambda,\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1}\right).$
We cast the following assumption on the parameter estimation error.
Assumption 2.
Suppose (^$(\widehat{}\mathbf{\Theta}_{1},\widehat{}\mathbf{\Theta}_{2})$ is independent of $\left(\mathbf{X}_{0},\mathbf{X}_{1}\right)$, and there exists $\gamma\in[0,0.04)$ such that for $\forall i\in[n]$ and $\forall\lambda\in\Lambda\cup\left\{\infty\right\}$,
$\mathbb{P}\left(\widetilde{Z}^{\lambda}_{0,i}\neq\bar{Z}_{0,i}^{\lambda}\right)\leq\gamma\cdot\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq Z_{0,i}^{*}\right).$
Remark 4.
To facilitate the analysis, we assume independence of (^$(\widehat{}\mathbf{\Theta}_{1},\widehat{}\mathbf{\Theta}_{2})$ and $\left(\mathbf{X}_{0},\mathbf{X}_{1}\right)$ in Assumption 2. In cases where the parameters are known (e.g., warm-up example in Section 2), this assumption holds with $\gamma=0$. In cases where the parameters need to be estimated, such independence can be achieved by either leveraging additional independent source data (e.g., in the context of item parameter calibration in LCMs and related latent variable models), or utilizing a sample splitting and label alignment trick on $\left(\mathbf{X}_{0},\mathbf{X}_{1}\right)$.
Under Assumption 2, we have the following observations:
  • (~$\left(\widetilde{}\mathbf{Z}^{\lambda}_{0},\widetilde{}\mathbf{Z}^{\lambda}_{1}\right)$ is coupled with (¯$\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\lambda}_{1}\right)$ through random seeds $\mathbf{E}_{0}$ and $\mathbf{E}_{1}$.
[…]
The following lemma indicates the validity of using $\widehat{\psi}$ in Algorithm 3 as our estimator for $\psi$, whose proof is given in Appendix D.3.
Lemma 3.
Suppose Assumption 1 and 2 hold. For the $\widehat{\psi}$ defined in Algorithm 3, with probability at least $1-\zeta$ we have
$\xi_{\widehat{\psi}}\leq C_{1}\left[\frac{\log\left(1/\zeta\right)}{n}+\gamma\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right)\right],$
and
$\widehat{\psi}\left(\lambda\right)\leq C_{2}\left[\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)+\frac{\log\left(1/\zeta\right)}{n}+\gamma\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right)\right],$
for some universal constants $C_{1},C_{2}>0$.
Combining Lemma 2 and Lemma 3, we immediately obtain the following result for our general transfer clustering procedure:
Theorem 5.
Suppose Assumption 1 and 2 hold, and there is a deterministic non-decreasing function $\phi$ on $\Lambda$ such that $\mathbb{P}\left({\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)\leq\phi\left(\lambda\right),\forall\lambda\in\Lambda\right)\geq 1-\zeta/2$. As a consequence, we have with probability at least $1-\zeta$ that
${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq C_{0}\left[\min_{\lambda\in\Lambda}\left\{\phi\left(\lambda\right)+\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)\right\}+\gamma\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right)+\frac{\log\left(1/\zeta\right)}{n}\right].$
for some universal constant $C_{0}>0$.
In Theorem 5, $\phi\left(\lambda\right)$ and $\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ can be viewed as the bias and the variance terms, respectively. The latter acts as a proxy for $\psi\left(\lambda\right)$. The additional term $\gamma\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right)$ accounts for the error introduced by the parameter estimation, which, together with $n^{-1}\log\left(1/\zeta\right)$, is typically negligible. In light of this, our adaptive method achieves the bias-variance trade-off over the grid $\Lambda$.
4.2 Two-component symmetric $d$-dimensional GMM
We now illustrate our general result using two-component symmetric multivariate GMM. Similar to the univariate case, let $\{Z_{0,i}^{*}\}_{i=1}^{n}$ and $\{Z_{1,i}^{*}\}_{i=1}^{n}$ be i.i.d. Rademacher latent variables such that $\mathbb{P}\big{(}Z_{0,i}^{*}\neq Z_{0,i}^{*}\big{)}\leq\varepsilon$. In addition, we have $\mathbf{X}_{0,i}|Z_{0,i}^{*}\sim N_{d}(Z_{0,i}^{*}{\bm{\mu}},\sigma^{2}\mathbf{I}_{d})$ and $\mathbf{X}_{1,i}|Z_{1,i}^{*}\sim N_{d}(Z_{0,i}^{*}{\bm{\mu}},\sigma^{2}\mathbf{I}_{d})$ for some ${\bm{\mu}}\in\mathbb{R}^{d}$ and $\sigma>0$, where $d\geq 1$. We investigate the performance of our method for this multivariate GMM with unknown $\left({\bm{\mu}},\sigma,\varepsilon\right)$ in the regime $\left\|{\bm{\mu}}\right\|/\sigma\rightarrow\infty$ and $\left\|{\bm{\mu}}\right\|/\sigma<\sqrt{2\log n}$. Suppose we have any estimator $\left(\widehat{\bm{\mu}},\widehat{\sigma}\right)$ that satisfies the following assumption:
Assumption 3.
Suppose $\left(\widehat{\bm{\mu}},\widehat{\sigma}\right)$ is independent of $\left(\mathbf{X}_{0},\mathbf{X}_{1}\right)$, and there exists an event ${\cal E}_{\sf par}$ with $\mathbb{P}\left({\cal E}_{\sf par}\right)\geq 1-n^{-C_{\sf par}}$ such that on ${\cal E}_{\sf par}$,
$\left|\frac{\min_{\theta\in\left\{\pm 1\right\}}\left\|\widehat{\bm{\mu}}-\theta{\bm{\mu}}\right\|}{\left\|{\bm{\mu}}\right\|}-1\right|+\left|\frac{\widehat{\sigma}}{\sigma}-1\right|\leq\frac{\delta_{n}}{4},$ (10)
for some $\delta_{n}=o(1)$ and $C_{\sf par}\geq 3$.
Applying TC to the two-component symmetric multivariate GMM for any $\lambda>0$ leads to
$\left(\widehat{Z}_{0,i}^{\lambda},\widehat{Z}_{1,i}^{\lambda}\right)$ $=\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\left\langle\widehat{\bm{\mu}},u\mathbf{X}_{0,i}+v\mathbf{X}_{1,i}\right\rangle+\lambda\widehat{\sigma}^{2}\mathbb{I}\left(u\neq v\right)\right\},\quad\forall i\in[n].$
For any grid set a grid set $\Lambda\subset[0,\infty)$, the following theorem provides an error bound for ^$\widehat{}\mathbf{Z}^{\widehat{\lambda}}_{0}=\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,\widehat{\bm{\mu}},\widehat{\sigma}\right)$, whose proof can be found in Appendix B.5.
Theorem 6.
Let $\zeta\in(0,1)$. Suppose Assumption 3 holds with $\delta_{n}\leq c_{0}\left(\log n\right)^{-3}$ and assume $B\geq C_{0}\zeta^{-2}\log\left(M\zeta^{-1}\right)$ for some universal constants $c_{0},C_{0}>0$. There exists some universal constant $C_{1}>0$ determined by $c_{0}$ and $C_{0}$ such that with probability at least $1-\zeta$,
(^$\ell\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{1}\min_{\lambda\in\Lambda}\left\{\mathcal{M}\left(\frac{\left\|{\bm{\mu}}\right\|}{\sigma},\varepsilon,\lambda\right)+\frac{\log\left(M\zeta^{-1}\right)}{n}\right\}.$
Analogous to Corollary 1, we have the following result certifying the adaptivity of ATC, whose proof is almost the same as that of Corollary 1 and hence omitted.
Corollary 2.
Suppose the conditions of Theorem 6 hold. Let $M=\lceil\log^{2}n\rceil$ and $\lambda_{j}=jM^{-1}\log n$ for $j=0,\cdots,M$, then there exists some universal constant $C_{2}>0$ determined by $C_{1}$ such that with probability at least $1-\zeta$,
(ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{2}\left[\mathcal{M}\left(\frac{\left\|{\bm{\mu}}\right\|}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right)+\frac{\log\left(M\zeta^{-1}\right)}{n}\right].$
Moreover, if ${\left\|{\bm{\mu}}\right\|}^{2}/\sigma^{2}<\left(1-c\right)\log n$ for some constant $c\in(0,1)$, then there exists some universal constant $C_{3}>0$ determined by $C_{2}$ such that with probability at least $1-n^{-10}$,
(ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{3}\mathcal{M}\left(\frac{\left\|{\bm{\mu}}\right\|}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right).$ (11)
Remark 5.
Our results can also be extended to the two-component symmetric Gaussian mixture model with unequal SNR, as considered in [19]. While the analysis follows a similar argument with more tedious calculations, the resulting expressions are more complex to interpret, hence we omit them for the sake of brevity.
5 Simulation Studies
In this section, we conduct numerical simulations to validate the effectiveness of our proposed estimator ^$\widehat{}\mathbf{Z}$ compared to ^$\widehat{}\mathbf{Z}^{\textsf{ITL}}$ and ^$\widehat{}\mathbf{Z}^{\textsf{DP}}$. Our code for reproducing results in this section is available at https://github.com/ZhongyuanLyu/ATC.
Given $\varepsilon\in[0,1/2]$, the random labels $\mathbf{Z}_{0}$ and $\mathbf{Z}_{1}$ are generated as follows:
  • Generate $Z_{0,i}\overset{i.i.d.}{\sim}\textsf{Unif}\left\{1,2\right\}$ for $i\in[n]$.
  • Generate $Z_{1,i}\in\left\{1,2\right\}$ by $\mathbb{P}\left(Z_{1,i}=Z_{0,i}\right)=1-\varepsilon$ and $\mathbb{P}\left(Z_{1,i}\neq Z_{0,i}\right)=\varepsilon$ for $i\in[n]$.
Given the label $\mathbf{Z}$ (either $\mathbf{Z}_{0}$ or $\mathbf{Z}_{1}$), we use the following procedure to generate our random samples:
  • (GMM) Randomly generate ${\bm{a}}\sim N_{d}\left(0,\mathbf{I}_{d}\right)$ and set ${\bm{\mu}}=\frac{{\bm{a}}}{\left\|{\bm{a}}\right\|}\cdot\upsilon$, then generate $\mathbf{X}_{i}\sim N_{d}\left({\bm{\mu}},\mathbf{I}_{d}\right)$ for $i\in[n]$.
  • (SBM) Set $\mathbf{B}=q\mathbf{1}_{K}\mathbf{1}_{K}^{\top}+(p-q)\mathbf{I}_{K}$ and generate $X_{ij}\sim\textsf{Bernoulli}(B_{Z_{i},Z_{j}})$ for $i,j\in[n]$.
  • (LCM) Set $\mathbf{\Theta}=[(0.5-\Delta)\mathbf{1}_{d},(0.5+\Delta)\mathbf{1}_{d}]$ and generate $X_{ij}\sim\textsf{Bernoulli}(\Theta_{j,Z_{i}})$ for $i\in[n]$ and $j\in[d]$.
In all simulations, we set $K=2$ and consider four different scenarios:
  • In the first setting (left panel in Figure 2), we use GMM as both target and source distribution with $(n,d,\upsilon)=(500,10,0.54)$.
  • In the second setting (right panel in Figure 2), we use SBM as target distribution and GMM as source distribution with $(n,d,p,q,\upsilon)=(300,10,0.4,0.3,0.79)$.
  • In the third setting (left panel in Figure 3), we use GMM as target distribution and SBM as source distribution with $(n,d,p,q,\upsilon)=(300,10,0.5,0.3,0.75)$.
  • In the fourth setting (right panel in Figure 3), we use LCM as target distribution and GMM as source distribution with $(n,d,\Delta,\upsilon)=(200,10,0.1,0.76)$.
  • In the fifth setting (Figure 4), we use LCM as both target and source distribution with $(n,d,\Delta,\upsilon)=(200,10,0.1,0.76)$.
In each setting, we fix $\upsilon$, $(p,q)$, or $\Delta$ (representing the signal-to-noise ratio in GMM, SBM and LCM, respectively), hence $\varepsilon$ controls the signal-to-noise ratio in transfer learning, analogous to $\alpha$ defined in Section 2.2. We apply our method with different choices of quantile parameter $1-\zeta\in\left\{0.8,0.9,0.95,0.99\right\}$. The numerical results are consistent with our theoretical findings. As shown in Figure 2-4, our estimator ^$\widehat{}\mathbf{Z}$ (ATC) outperforms, or at least matches, ^$\widehat{}\mathbf{Z}^{\textsf{ITL}}$ and ^$\widehat{}\mathbf{Z}^{\textsf{DP}}$ (denoted by ITL and DP, respectively), which indicates the benefit of our proposed method. In addition, it can be seen that our method is robust to the choice of $\zeta$.
Figure 2: Clustering error v.s. discrepancy $\varepsilon$. (Left) Target: GMM; Source: GMM. Parameters are set to be $(n,K,d,\upsilon)=(500,2,10,0.54)$. (Right) Target: SBM; Source: GMM. Parameters are set to be $(n,K,d,p,q,\upsilon)=(300,2,10,0.4,0.3,0.79)$. Each error point represents the average of 50 replications.
Figure 3: Clustering error v.s. discrepancy $\varepsilon$. (Left) Target: GMM; Source: SBM. Parameters are set to be $(n,K,p,q,d,\upsilon)=(300,2,0.5,0.3,10,0.75)$. (Right) Target: LCM; Source: GMM. Parameters are set to be $(n,K,d,\Delta,\upsilon)=(200,2,10,0.1,0.76)$. Each error point represents the average of 50 replications.
Figure 4: Clustering error v.s. discrepancy $\varepsilon$. Target: LCM; Source: LCM. Parameters are set to be $(n,K,d,\Delta)=(300,2,15,0.1)$. Each error point represents the average of 50 replications.
6 Real data applications
In this section, we evaluate the performance of ATC on three real-world datasets, including the Lazega Lawyers Network data, TIMSS 2019 educational assessment data, and Business Relation Network data. Our code for reproducing results in this section is available at https://github.com/ZhongyuanLyu/ATC.
6.1 Lazega Lawyers Network
The Lazega Lawyers Network originates from a network study of corporate law partnership conducted in a US corporate law firm in 1988-1991 [39]. The dataset includes a strong-coworker network among the 71 attorneys of this firm. For each attorney, we selected years with the firm as the covariate and status (either partner or associate) as the true label. The misclassification error is $0.151$ when using only the covariate and $0.394$ when using only the network. This suggests that the covariate largely reflects the clustering structure, while the network provides limited information regarding the clustering of attorney status. We treat the covariate as the target data and the network as the source data and apply ATC with $1-\zeta\in\left\{0.8,0.9,0.95\right\}$.
In Figure 5, the ground truth line is the misclassification error of TC varying with $\lambda$. Other lines are the estimated error, i.e., $\widehat{\phi}(\lambda)+\widehat{\psi}(\lambda)$ based on ATC without using the true labels. As shown in Figure 5, ATC can adaptively choose the smallest $\lambda$ that minimizes the true misclassification error across all choices of $\zeta$, which is $0.076$. This indicates that our method effectively integrates information from both the covariate and the network, resulting in improved clustering performance.
Figure 5: (Clustering error v.s. penalty $\lambda$ for Lawyers Network) Target: years with the firm (GMM); Source: strong-coworker network (SBM).
6.2 Trends in International Mathematics and Science Study Data
We investigate the performance of our method on TIMSS 2019 dataset (Trends in International Mathematics and Science Study) [^3] The data is publicly available at https://timssandpirls.bc.edu/timss2019. We used the student achievement data and context data at eighth grades in Singapore, which has highest average score across all participating countries. We preprocess the data by selecting $n=259$ students that have access to a common set of items related to science and mathematics. We collect the feature that represents to what extent the student would agree to the statement “My teacher tells me I am good at science”, categorized by $1$: Agree a lot; $2$: Agree a little; $3$: Disagree a little; $4$: Disagree a lot. We further label $\left\{1\right\}$ as $1$ and $\left\{2,3,4\right\}$ as 0, indicating whether the student is good at science or not as our ground truth. The target data comprised students’ responses to $d=17$ science-related items, while the source data included their responses to $d=17$ mathematics-related items. We use the latent class model with two classes for both the source data and the target data.
Left panel in Figure 6 presents the performance of ATC for $1-\zeta\in\left\{0.8,0.9,0.95,0.99\right\}$. The misclassification error was $0.371$ using the target data. In contrast, ATC achieves a lower misclassification error of $0.347$ across all choices of $\zeta$.
To further quantify the uncertainty of ATC, we present a violin plot (right panel in Figure 6) showing the misclassification errors based on $N_{b}=100$ bootstrap samples (with replacement) from the original $n=259$ samples. The plot clearly demonstrates that our method exhibits greater robustness compared to DP and ITL.
To facilitate interpretation, we present a heatmap with the the estimated conditional probabilities of giving the correct response to each item for each latent class, as shown in Figure 7. A closer inspection of the items at the bottom of the figure reveals that these are sub-questions of a single overarching question: “Components Julia must use to build a circuit.” The design of these items appears to have lower discriminative power, as students in both classes tend to answer them correctly.
Figure 6: (Left) Clustering error v.s. penalty $\lambda$ for TIMSS data. Target: responses to science-related items (LCM); Source: responses to math-related items (LCM). (Right) Clustering error via resampling for TIMSS data. Violin plot is based on $N_{b}=100$ bootstrap (sample with replacement) of original $n=259$ samples.
Figure 7: Item parameters estimated by ATC in the latent class model for the TIMSS data.
Figure 8: (Clustering error v.s. penalty $\lambda$ for Business Relation Network) Target: supplier network (SBM); Source: standardized closing prices (GMM).
Type DP Target Oracle ATC(0.8) ATC(0.9) ATC(0.95) ATC(0.99)
Misclassification 0.551 0.734 0.532 0.535 0.535 0.535 0.535
ARI 0.129 0.003 0.138 0.134 0.134 0.134 0.134
Table 2: Error rates for business relation network
6.3 Business Relation Network
We then test our method on the Business Relation Network with contextual information, publicly available at https://data.world/datasyndrome/relato-business-graph-database. We use the data pre-processed by [62], from which we consider the supplier network among $n=312$ companies as the target data, and covariate vectors of dimension $d=102$ being the standardized closing prices of these companies in the stock market from 2021-01-01 to 2021-06-01 as the source data. The companies we consider here belong to $K=5$ sectors, including “health”, “information technology”, “financials”, “consumer discretionary”, and “industrials”, which are treated as their true community labels.
Figure 8 shows the performance of ATC with different $1-\zeta\in\left\{0.8,0.9,0.95,0.99\right\}$. It is suggested that we shall choose a large $\lambda$, which has a similar performance as data pooling. The misclassification error is $0.734$ by using the target data, which indicates the extremely low SNR in this dataset. By our adaptive choice of $\lambda$, we are able to achieve a lower misclassification error of $0.535$ for all choices of $\zeta$. It’s worth mentioning that the best misclassification error reported in [62] is $0.599$, which utilizes $5$ layers of network.
7 Disscussion
In this paper, we have developed a unified transfer learning framework for clustering in the presence of unknown discrepancy between latent labels of target and source data, applicable to a wide range of statistical models. The proposed adaptive transfer clustering (ATC) algorithm effectively balances the bias-variance trade-off and its theoretical optimality under two-component symmetric GMMs is well established. Several future directions are worth exploring. It would be of great interest to extend the analysis in Theorem 5 to LCMs and SBMs, which requires more technical efforts. Moreover, our framework can be adapted to other unsupervised learning tasks beyond clustering. It would be intriguing to examine its empirical and theoretical performance in continuous parameter estimation in models such as mixed membership models and continuous latent factor models.
Acknowledgement
Kaizheng Wang’s research is supported by an NSF grant DMS-2210907 and a startup grant at Columbia University. Yuqi Gu’s research is supported by an NSF grant DMS-2210796 and startup funding from the Department of Statistics at Columbia University.
Appendix
The appendix provides all the proofs of theoretical results.
Appendix A Auxiliary lemmas
Lemma 4.
Let $\left\{X_{0,i}\right\}_{i=1}^{n}$ be i.i.d. Ber$\left(p\right)$. For any $\vartheta\in(0,1],\delta\in(0,1)$ we have
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{0,i}\leq\left(1+\vartheta\right)p+\frac{7\log\left(1/\delta\right)}{6\vartheta n}\right)\geq 1-\delta,$
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{0,i}\geq\left(1-\vartheta\right)p-\frac{7\log\left(1/\delta\right)}{6\vartheta n}\right)\geq 1-\delta.$
Proof: see Section D.4.
Lemma 5.
For any $x>0$, we have
$\frac{\sqrt{2/\pi}}{x+\sqrt{x^{2}+4}}\exp\left(-\frac{x^{2}}{2}\right)\leq\Phi\left(-x\right)\leq\frac{\sqrt{2/\pi}}{x+\sqrt{x^{2}+8/\pi}}\exp\left(-\frac{x^{2}}{2}\right).$
Lemma 6.
For any $x\in\mathbb{R}$ and $\Delta>0$, we have
$\Phi\left(x+\Delta\right)-\Phi\left(x\right)\leq\frac{\Delta}{\sqrt{2\pi}}\exp\left(-\frac{\min\left\{x^{2},\left(x+\Delta\right)^{2}\right\}}{2}\right),$
and
$\Phi\left(x+\Delta\right)-\Phi\left(x\right)\geq\frac{\Delta}{\sqrt{2\pi}}\exp\left(-\frac{\max\left\{x^{2},\left(x+\Delta\right)^{2}\right\}}{2}\right).$
Proof: The result immediately follows by definition of $\Phi$.
We also present the following assumption and lemma which are needed in our main proofs.
Assumption 4.
Consider $\left(\mathbf{X}_{0},\mathbf{X}_{1}\right)$ which follows the two-component symmetric univariate GMM model with parameter $\left(\mu,\sigma,\varepsilon\right)$. Suppose $\left(\widehat{\mu},\widehat{\sigma}\right)$ is independent of $\left(\mathbf{X}_{0},\mathbf{X}_{1}\right)$, and there exists an event ${\cal E}_{\sf par}$ with $\mathbb{P}\left({\cal E}_{\sf par}\right)\geq 1-n^{-C_{\sf par}}$ such that on ${\cal E}_{\sf par}$,
$\max\left\{\frac{\left|{\widehat{\mu}}/{\widehat{\sigma}}-{\mu}/{\sigma}\right|}{\mu/\sigma},\frac{\left|\widehat{\sigma}/{\widehat{\mu}}-\sigma/{\mu}\right|}{\sigma/{\mu}}\right\}\leq\delta_{n},$ (12)
for some $\delta_{n}=o(1)$ and $C_{\sf par}\geq 3$.
Lemma 7.
Suppose Assumption 4 holds. For each $i\in[n]$,
$\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}\right)\geq$ $\max\left\{\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)\left(1-2\Phi\left(-\frac{\lambda}{\sqrt{2}\mu/\sigma}\right)\right),\left[\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}\right\}$
$+\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right).$
Proof: see Section D.5.
Appendix B Proofs of Main Results
Throughout the proof, we write $\mathbb{P}_{a,b}\left(\cdot\right):=\mathbb{P}\left(\cdot\mid(Z^{*}_{0,i},Z^{*}_{1,i})=(a,b)\right)$ for brevity.
B.1 Proof of Thereom 1
Define
$L(u,v)=\mu\left(uX_{0,i}+vX_{1,i}\right)+\lambda\sigma^{2}\mathbb{I}\left(u\neq v\right),\qquad u,v\in\{\pm 1\}.$
If $\check{Z}^{\lambda}_{0,i}=-1$, then
$L(-1,1)$ $\leq\min\{L(1,1),L(1,-1),L(-1,-1)\}\quad$
$\text{or}\qquad L(-1,-1)$ $\leq\min\{L(1,1),L(1,-1),L(-1,1)\}$
should happen. Let $K=\frac{\lambda\sigma^{2}}{2\mu}$. The aforementioned two events translate to
$X_{0,i}\leq X_{1,i},\quad X_{0,i}\leq-K,\quad X_{1,i}\geq K$ (13)
and
$X_{0,i}+X_{1,i}\leq 0,\quad X_{0,i}\leq K,\quad X_{1,i}\leq K,$ (14)
respectively. We start with the case $(Z^{*}_{0,i},Z^{*}_{1,i})=(1,1)$. Note that
[…]
and
[…]
Hence,
$\mathbb{P}_{1,1}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i})\leq\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+\Phi\bigg{(}\frac{-K-\mu}{\sigma}\bigg{)}\Phi\bigg{(}\frac{-K+\mu}{\sigma}\bigg{)}.$ (15)
We then consider the case $(Z^{*}_{0,i},Z^{*}_{1,i})=(1,-1)$. Note that
[…]
and
[…]
Thus we can conclude that
$\mathbb{P}_{1,-1}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i})\leq\Phi\left(\frac{K-\mu}{\sigma}\right)+\Phi\left(\frac{-K-\mu}{\sigma}\right).$ (16)
By symmetry, the right-hand sides of (15) and (16) can upper bound $\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}=Z^{*}_{1,i})$ and $\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}\neq Z^{*}_{1,i})$, respectively. Therefore,
$\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i})$
$=\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}=Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}=Z^{*}_{1,i})+\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}\neq Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i})$
$\leq\bigg{[}\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+\Phi\bigg{(}\frac{-K-\mu}{\sigma}\bigg{)}\Phi\bigg{(}\frac{-K+\mu}{\sigma}\bigg{)}\bigg{]}\cdot[1-\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i})]$
$\quad+\bigg{[}\Phi\left(\frac{K-\mu}{\sigma}\right)+\Phi\left(\frac{-K-\mu}{\sigma}\right)\bigg{]}\cdot\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i})$
$=\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+\Phi\bigg{(}\frac{-K-\mu}{\sigma}\bigg{)}\Phi\bigg{(}\frac{-K+\mu}{\sigma}\bigg{)}$
$\quad+\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i})\cdot\bigg{\{}\Phi\bigg{(}\frac{K-\mu}{\sigma}\bigg{)}\bigg{[}1+\Phi\bigg{(}\frac{-K-\mu}{\sigma}\bigg{)}\bigg{]}-\Phi\bigg{(}-\frac{\sqrt{2}\mu}{\sigma}\bigg{)}\bigg{\}}.$
We conclude that
$\mathbb{P}(\check{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i})\leq\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+\Phi\bigg{(}-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\bigg{)}\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)+2\varepsilon\Phi\bigg{(}-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\bigg{)}.$
B.2 Proof of Theorem 2
Without loss of generality we assume $\sigma=1$, otherwise we can always proceed our analysis on standardized data $\left\{X_{0,i}/\sigma,i\in[n]\right\}$. We take $\lambda^{*}=\log\left(1/\varepsilon\right)$. By Proposition 1, we obtain that
$\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}\right)$
$\leq\left(1-\varepsilon\right)\left[\Phi\left(-\sqrt{2}\mu\right)+\Phi\left(-\left(\frac{\log\left(1/\varepsilon\right)}{2\mu}-\mu\right)\right)\Phi\left(-\mu-\frac{\log\left(1/\varepsilon\right)}{2\mu}\right)\right]$
$+\varepsilon\left[2\Phi\left(-\left(\mu-\frac{\log\left(1/\varepsilon\right)}{2\mu}\right)\right)\wedge 1\right].$
We first consider the case when $10\leq\alpha<\sqrt{2}-$ 1, for which we obtain that
$\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}\right)$
$\leq\Phi\left(-\sqrt{2}\mu\right)+\Phi\left(-\mu-\frac{\log\left(1/\varepsilon\right)}{2\mu}\right)+2\exp\left(-2\mu^{2}\cdot\alpha\right)\Phi\left(-\left(\mu-\frac{\log\left(1/\varepsilon\right)}{2\mu}\right)\right)$
$\leq\frac{1}{2\sqrt{\pi}\mu}\exp\left(-\mu^{2}\right)+\frac{1}{\sqrt{2\pi}(1+\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}\left(1+\alpha\right)^{2}\right)$
$+\frac{2}{\sqrt{2\pi}(1-\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}(1+\alpha)^{2}\right)$
$\leq\frac{1}{2\sqrt{\pi}\mu}\exp\left(-\mu^{2}\right)+\frac{3}{\sqrt{2\pi}(1-\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}(1+\alpha)^{2}\right)$
$\leq\frac{C_{1}}{\mu}\exp\left(-\frac{\mu^{2}}{2}(1+\alpha)^{2}\right),$
for some universal constant $C_{1}>0$. where the last inequality holds since $(1+\alpha)^{2}\leq 2$.
It remains to consider $\alpha>\sqrt{2}-1$. If $\sqrt{2}-1<\alpha<1/2$, then
$\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}\right)$
$\leq\Phi\left(-\sqrt{2}\mu\right)+\Phi\left(-\mu-\frac{\log\left(1/\varepsilon\right)}{2\mu}\right)+2\exp\left(-2\mu^{2}\cdot\alpha\right)\Phi\left(-\left(\mu-\frac{\log\left(1/\varepsilon\right)}{2\mu}\right)\right)$
$\leq\frac{1}{2\sqrt{\pi}\mu}\exp\left(-\mu^{2}\right)+\frac{1}{\sqrt{2\pi}(1+\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}\left(1+\alpha\right)^{2}\right)$
$+\frac{2}{\sqrt{2\pi}(1-\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}(1+\alpha)^{2}\right)$
$\leq\frac{1}{2\sqrt{\pi}\mu}\exp\left(-\mu^{2}\right)+\frac{3}{\sqrt{2\pi}(1-\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}(1+\alpha)^{2}\right)$
$\leq\frac{\widetilde{C}_{2}}{\mu}\exp\left(-{\mu^{2}}\right).$
for some universal constant $\widetilde{C}_{2}>0$. If $\alpha\geq 1/2$ then
$\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}\right)$
$\leq\Phi\left(-\sqrt{2}\mu\right)+\Phi\left(-\mu-\frac{\log\left(1/\varepsilon\right)}{2\mu}\right)+\exp\left(-2\mu^{2}\cdot\alpha\right)$
$\leq\frac{1}{2\sqrt{\pi}\mu}\exp\left(-\mu^{2}\right)+\frac{1}{\sqrt{2\pi}(1+\alpha)\mu}\exp\left(-\frac{\mu^{2}}{2}\left(1+\alpha\right)^{2}\right)+\exp\left(-2\mu^{2}\cdot\alpha\right)$
$\leq\frac{1}{\sqrt{\pi}\mu}\exp\left(-{\mu^{2}}\right)+\exp\left(-\mu^{2}\cdot 2\alpha\right).$
The proof is completed by collecting all bounds for three cases.
B.3 Proof of Theorem 3
Fix $i\in[n]$, denote the posterior log-likelihood function for $(Z^{*}_{0,i},Z^{*}_{1,i})$ as
$g\left(z_{0},z_{1}\mid X_{0},X_{1}\right)=\log\mathbb{P}\left(Z^{*}_{0,i}=z_{0},Z^{*}_{1,i}=z_{1}\mid X_{0,i}=x_{0},X_{1,i}=x_{1}\right).$
It is easy to see that
$g\left(z_{0},z_{1}\mid x_{0},x_{1}\right)$ $=\frac{\mu}{\sigma^{2}}\left(z_{0}x_{0}+z_{1}x_{1}\right)-\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\mathbb{I}\left(z_{0}\neq z_{1}\right)+C\left(x_{0},x_{1}\right)$
where $C\left(x_{0},x_{1}\right)$ is a function independent of $(z_{0},z_{1})$. For any $\mathbf{Z},\mathbf{Z}^{\prime}\in\left\{\pm 1\right\}^{n}$, define $s(\mathbf{Z},\mathbf{Z}^{\prime})=\operatorname*{arg\,min}_{u\in\left\{\pm 1\right\}}\sum_{i=1}^{n}\mathbb{I}\left(Z_{i}\neq uZ_{i}^{\prime}\right)$. Notice that for any estimator (^$(\widehat{}\mathbf{Z}_{0},\widehat{}\mathbf{Z}_{1})$ and any $i\in[n]$, we have
$\mathbb{P}\left(\left(\widehat{Z}_{0,i},\widehat{Z}_{1,i}\right)\neq\left(Z^{*}_{0,i},Z^{*}_{1,i}\right)\right)$
$\geq\mathbb{P}\left(g\left(Z^{*}_{0,i},Z^{*}_{1,i}\mid X_{0,i},X_{1,i}\right)<\max_{\left(z_{0},z_{1}\right)\neq\left(Z^{*}_{0,i},Z^{*}_{1,i}\right)}g\left(z_{0},z_{1}\mid X_{0,i},X_{1,i}\right)\right).$ (17)
where the inequality holds due to the Bayes optimal estimator under 00-$1$ loss is given by the posterior mode (MAP). It suffices to lower bound the event inside the RHS of (B.3), denoted by ${\cal E}_{\sf lb}$.
Define $K=\frac{\sigma^{2}}{2\mu}\log\left(\frac{1-\varepsilon}{\varepsilon}\right)$. When $\left(Z^{*}_{0,i},Z^{*}_{1,i}\right)=\left(1,1\right)$, ${\cal E}_{\sf lb}$ is equivalent to
$\left\{X_{0,i}<-K\right\}\cup\left\{X_{1,i}<-K\right\}\cup\left\{X_{0,i}+X_{1,i}<0\right\}.$
Hence we have
$\mathbb{P}_{1,1}\left({\cal E}_{\sf lb}\right)$
$=\mathbb{P}_{1,1}\left(X_{0,i}<-K\right)+\mathbb{P}_{1,1}\left(X_{1,i}<-K\right)+\mathbb{P}_{1,1}\left(X_{0,i}+X_{1,i}<0\right)$
$\quad-\mathbb{P}_{1,1}\left(\left\{X_{0,i}<-K\right\}\bigcap\left\{X_{0,i}+X_{1,i}<0\right\}\right)-\mathbb{P}_{1,1}\left(\left\{X_{1,i}<-K\right\}\bigcap\left\{X_{0,i}+X_{1,i}<0\right\}\right)$
$=\mathbb{P}_{1,1}\left(\left\{X_{0,i}<-K\right\}\bigcap\left\{X_{0,i}+X_{1,i}>0\right\}\right)+\mathbb{P}_{1,1}\left(\left\{X_{1,i}<-K\right\}\bigcap\left\{X_{0,i}+X_{1,i}>0\right\}\right)$
$\quad+\mathbb{P}_{1,1}\left(X_{0,i}+X_{1,i}<0\right)$
$=\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+2\mathbb{P}_{1,1}\left(\left\{X_{0,i}<-K\right\}\bigcap\left\{X_{0,i}+X_{1,i}>0\right\}\right).$
[…]
$\mathbb{P}_{1,1}\left(\left\{X_{0,i}<-K\right\}\cap\left\{X_{0,i}+X_{1,i}>0\right\}\right)\geq\frac{1}{2}\mathbb{P}_{1,1}\left(X_{0,i}<-K\right)\mathbb{P}_{1,1}\left(X_{1,i}>K\right).$ (18)
To see it, let ${\cal S}_{0}:=\left\{\left(x,y\right)\in\mathbb{R}^{2}:x<-K,x+y>0\right\}$, ${\cal S}_{1}:=\left\{\left(x,y\right)\in\mathbb{R}^{2}:y>K,x+y<0\right\}$ and $p\left(x\right):=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{\left(x-\mu\right)^{2}}{2\sigma^{2}}\right)$. Define $I_{0}:=\int_{{\cal S}_{0}}p(x)p(y)\mathrm{d}x\mathrm{d}y$ and $I_{1}:=\int_{{\cal S}_{1}}p(x)p(y)\mathrm{d}x\mathrm{d}y$. By definition we have $I_{0}=\mathbb{P}_{1,1}\left(\left\{X_{0,i}<-K\right\}\bigcap\left\{X_{0,i}+X_{1,i}>0\right\}\right)$ and
$I_{0}+I_{1}=\mathbb{P}_{1,1}\left(X_{0,i}<-K\right)\mathbb{P}_{1,1}\left(X_{1,i}>K\right).$
To show (18), it boils down to show $I_{0}\geq I_{1}$. Notice that
$I_{1}$ $=\int_{{\cal S}_{1}}p(x)p(y)\mathrm{d}x\mathrm{d}y=\int_{{\cal S}_{0}}p(-y)p(-x)\mathrm{d}x\mathrm{d}y$ (By change of variable)
$=\int_{{\cal S}_{0}}\frac{1}{2\pi\sigma^{2}}\exp\left(-\frac{\left(x+\mu\right)^{2}}{2\sigma^{2}}\right)\exp\left(-\frac{\left(y+\mu\right)^{2}}{2\sigma^{2}}\right)\mathrm{d}x\mathrm{d}y$
$=\int_{{\cal S}_{0}}p(x)p(y)e^{-\frac{\mu}{\sigma^{2}}\left(x+y\right)}\mathrm{d}x\mathrm{d}y$
[…]
Hence we can conclude that (18) holds. Then,
$\mathbb{P}_{1,1}\left({\cal E}_{\sf lb}\right)\geq\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+\Phi\left(\frac{-\mu-K}{\sigma}\right)\Phi\left(\frac{\mu-K}{\sigma}\right).$ (19)
When $\left(Z^{*}_{0,i},Z^{*}_{1,i}\right)=\left(1,-1\right)$, ${\cal E}_{\sf lb}$ is equivalent to
$\left\{X_{0,i}<K\right\}\cup\left\{X_{1,i}>-K\right\}\cup\left\{X_{0,i}<X_{1,i}\right\}.$
Hence we have
$\mathbb{P}_{1,-1}\left({\cal E}_{\sf lb}\right)=1-\mathbb{P}_{1,-1}\left(\left\{X_{0,i}>K\right\}\cap\left\{X_{1,i}<-K\right\}\cap\left\{X_{0,i}>X_{1,i}\right\}\right)$
$=1-\mathbb{P}_{1,-1}\left(X_{0,i}>K\right)\mathbb{P}_{1,-1}\left(X_{1,i}<-K\right)$
$=1-\left[\mathbb{P}_{1,-1}\left(X_{0,i}>K\right)\right]^{2}$
$=1-\Phi^{2}\bigg{(}\frac{\mu-K}{\sigma}\bigg{)}\geq\mathbb{P}_{1,-1}\left(X_{0,i}<K\right)=\Phi\left(\frac{K-\mu}{\sigma}\right),$ (20)
[…]
$\mathbb{P}\left(\left(\widehat{Z}_{0,i},\widehat{Z}_{1,i}\right)\neq\left(Z^{*}_{0,i},Z^{*}_{1,i}\right)\right)$
$\geq(1-\varepsilon)\left[\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+\Phi\left(\frac{-\mu-K}{\sigma}\right)\Phi\left(\frac{\mu-K}{\sigma}\right)\right]+\varepsilon\Phi\left(\frac{K-\mu}{\sigma}\right).$
We finish the proof by using the fact that
$\mathbb{P}\left((\widehat{Z}_{0,i},\widehat{Z}_{1,i})\neq\left(Z^{*}_{0,i},Z^{*}_{1,i}\right)\right)\leq\mathbb{P}\left(\widehat{Z}_{0,i}\neq Z^{*}_{0,i}\right)+\mathbb{P}\left(\widehat{Z}_{1,i}\neq Z^{*}_{1,i}\right)=2\mathbb{P}\left(\widehat{Z}_{0,i}\neq Z^{*}_{0,i}\right).$
B.4 Proofs of Theorems 4 and 7
Note that Theorem 4 is a special case of Theorem 7, Assumption 4 holds with $\delta_{n}=0$. Hence it suffices to prove Theorem 7.
We will invoke Lemma 2. To that end, we first need to show that
((^$\mathbb{P}\left(\ell\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)\leq\phi\left(\lambda\right),\forall\lambda\in\Lambda\right)\geq 1-\zeta$
with some $\zeta\in(0,1)$ and a non-decreasing $\phi$. Notice that,
(^$\ell\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)\leq\ell\left(\check{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)+\ell\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\check{}\mathbf{Z}_{0}^{\lambda}\right),\quad\lambda\in\Lambda.$
We need the following two lemmas whose proofs are deferred to Section D.6 and Section D.7.
Lemma 8.
For each $i\in[n]$ and $\lambda\geq 0$, we have
$\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq\bar{Z}^{\lambda}_{0,i}\right)$ $\leq 4\min\left\{\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right),\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)\right\},$
$\mathbb{P}\left(\check{Z}_{0,i}^{\lambda}\neq\bar{Z}_{0,i}^{\lambda}\right)$ $\leq\varepsilon\left[2\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right)\wedge 1\right].$
Lemma 9.
Suppose Assumption 4 holds with $\delta_{n}=o\left(\left(\mu/\sigma\right)^{-2}\right)$, then for each $i\in[n]$ and $\lambda\geq 0$ we have
$\mathbb{P}\left(\widehat{Z}_{0,i}^{\lambda}\neq\check{Z}_{0,i}^{\lambda}\right)$ $\leq\frac{C\delta_{n}\lambda\left(1+\delta_{n}\lambda\right)}{\mu/\sigma}\left[\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)+\varepsilon\cdot\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right)\right]+n^{-C_{\sf par}}.$
where $C>0$ is an universal constant.
By Lemma 8 and Lemma 9, we get that
$\mathbb{E}\ell\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)\leq\mathbb{E}\ell\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\check{}\mathbf{Z}_{0}^{\lambda}\right)+\mathbb{E}\ell\left(\check{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)\leq 3\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right).$
provided that $\lambda\lesssim\log n$, $\delta_{n}=o\left(\left(\log n\right)^{-1}\right)$ due to $\mu/\sigma\lesssim\sqrt{\log n}$ and $\delta_{n}=o\left(\left(\frac{\mu/\sigma}{\log n}\right)^{2}\right)$, and $C_{\sf par}\geq 3$. Combined with Lemma 4, with probability at least $21-\zeta/$ 2 that for any $\lambda\in\Lambda$,
(^$\ell\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)\leq 4\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right)+\frac{7\log\left(2M/\zeta\right)}{2n}=:\phi\left(\lambda\right)$
It is readily seen that $\phi\left(\lambda\right)$ is non-decreasing. By Lemma 2, we obtain that with probability at least $21-\zeta/$ 2,
(^$\ell\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq 4\min_{\lambda\in\Lambda}\left\{\phi\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right\}+3\xi_{\widehat{\psi}}.$ (21)
Next we will invoke Lemma 3. To that end, we need the following lemma to verify Assumption 2 holds under Assumption 4, whose proof can be found in Section D.8.
Lemma 10.
Suppose Assumption 4 holds with $\delta_{n}=o\left(\frac{1}{\left(\mu/\sigma\right)^{3}}\wedge\left(\frac{\mu/\sigma}{\log n}\right)^{2}\right)$ and $\max_{\lambda\in\Lambda}\lambda\lesssim\log n$, then Assumption 2 is satisfied for two-component symmetric GMM with $\gamma=o(1)$.
By Lemma 3 and Lemma 10 with $\gamma=0.004$, we can conclude that with probability at least $21-\zeta/$ 2,
$\xi_{\widehat{\psi}}\leq\frac{2105\log\left(16/\zeta\right)}{n}+2\mathbb{E}\ell\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right),$
and
$\widehat{\psi}\left(\lambda\right)\leq 8\mathbb{E}\ell\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)+\frac{42\log\left(32/\zeta\right)}{n}+0.05\mathbb{E}\ell\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right).$
Combined with (21), we can conclude that with probability at least $1-\zeta$,
(^$\ell\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq C\min_{\lambda\in\Lambda}\left[\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right)+\mathbb{E}\ell\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)+\mathbb{E}\ell\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right)+\frac{\log\left(M/\zeta\right)}{n}\right].$
for some universal constant $C>0$. By using $\mathbb{E}\ell\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right)=\Phi\left(-\sqrt{2}\mu/\sigma\right)$ and Lemma 8, we obtain that with probability at least $1-\zeta$,
$\ell$ (^$\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ (22)
$\leq\widetilde{C}\min_{\lambda\in\Lambda}\left[\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)+\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right)+\frac{\log\left(M/\zeta\right)}{n}\right].$
On the other hand, by Lemma 4 we have
((¯$\mathbb{P}\left(\ell\left(\bar{}\mathbf{Z}_{0}^{\infty},\mathbf{Z}_{0}^{*}\right)\leq 2\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)+\frac{7\log\left(2/\zeta\right)}{6n}\right)\geq 1-\zeta/2.$ (23)
The proof is completed by using triangle inequality and combining (22) and (23).
B.5 Proof of Theorem 6
We first prove the result for $d=1$ using the original samples $\left(\mathbf{X}_{0},\mathbf{X}_{1}\right)$ with unknown $\left(\mu,\sigma,\varepsilon\right)$ under 4, as the proof for $d\geq 2$ can be reduced to univariate case.
The following theorem can be regarded as a generalized version of Theorem 4 fully adaptive to $\left(\mu,\sigma,\varepsilon\right)$ in the univariate case, whose proof can be found in Section B.4.
Theorem 7.
Let ^$\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}}:=\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,\widehat{\mu},\widehat{\sigma}\right)$. Suppose Assumption 4 holds with $\delta_{n}=o\left(\frac{1}{\left(\mu/\sigma\right)^{3}}\wedge\left(\frac{\mu/\sigma}{\log n}\right)^{2}\right)$. Let $\zeta\in(0,1)$, and assume that $B\geq C_{0}\zeta^{-2}\log\left(M\zeta^{-1}\right)$ for some universal constant $C_{0}>0$, then with probability at least $1-\zeta$, we have
(^$\ell\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{1}\min_{\lambda\in\Lambda}\left\{\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\lambda\right)+\frac{\log\left(M\zeta^{-1}\right)}{n}\right\}.$
for some universal constant $C_{1}>0$.
Now we turn to the proof for $d\geq 2$. By definition, for $\forall i\in[n]$ we have
$\left(\widehat{Z}_{0,i}^{\lambda},\widehat{Z}_{1,i}^{\lambda}\right)$ $=\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\left\langle\widehat{\bm{\mu}},u\mathbf{X}_{0,i}+v\mathbf{X}_{1,i}\right\rangle+\lambda\widehat{\sigma}^{2}\mathbb{I}\left(u\neq v\right)\right\}$ (24)
$=\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\left\|\widehat{\bm{\mu}}\right\|\left(\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}_{0,i}}{\left\|\widehat{\bm{\mu}}\right\|}+\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}_{1,i}}{\left\|\widehat{\bm{\mu}}\right\|}\right)+\lambda\widehat{\sigma}^{2}\mathbb{I}\left(u\neq v\right)\right\}.$
In view of (24), we can construct projected samples $\big{\{}\widetilde{X}_{0,i},\widetilde{X}_{1,i}\big{\}}_{i=1}^{n}$ using the estimated parameters $\left(\widehat{\bm{\mu}},\widehat{\sigma}\right)$:
$\widetilde{X}_{0,i}:=\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}_{0,i}}{\left\|\widehat{\bm{\mu}}\right\|},\quad\widetilde{X}_{1,i}:=\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}_{1,i}}{\left\|\widehat{\bm{\mu}}\right\|},\quad\forall i\in[n].$
We thus have $\textsf{TC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\lambda,{\widehat{\bm{\mu}}},\widehat{\sigma}\right)=\textsf{TC}\big{(}\widetilde{}\mathbf{X}_{0},\widetilde{}\mathbf{X}_{1};\lambda,\left\|\widehat{\bm{\mu}}\right\|,\widehat{\sigma}\big{)}$ for any $\lambda>0$. Note that we can view $\left(\left\|\widehat{\bm{\mu}}\right\|,\widehat{\sigma}\right)$ as fixed parameters, which serves as a good approximation of $\left(\left\|{\bm{\mu}}\right\|,\sigma\right)$ by 3. Denote $\mu:=\left\|{\bm{\mu}}\right\|$ and $\widetilde{\mu}:=\left\|\widehat{\bm{\mu}}\right\|^{-1}{\widehat{\bm{\mu}}^{\top}{\bm{\mu}}}$. Due to independence, we have
$\widetilde{X}_{0,i}|Z_{0,i}^{*}\sim N\big{(}Z_{0,i}^{*}\widetilde{\mu},\sigma^{2}\big{)},\quad\widetilde{X}_{1,i}|Z_{1,i}^{*}\sim N\big{(}Z_{1,i}^{*}\widetilde{\mu},\sigma^{2}\big{)},\quad\forall i\in[n].$
In addition, let
$\widetilde{X}^{b}_{0,i}:=\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}^{b}_{0,i}}{\left\|\widehat{\bm{\mu}}\right\|},\quad\widetilde{X}_{1,i}:=\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}^{b}_{1,i}}{\left\|\widehat{\bm{\mu}}\right\|},\quad\forall i\in[n],\quad\forall b\in[B].$
We thus have
$\widetilde{X}_{0,i}^{b}|Z_{0,i}^{*}\sim N\big{(}Z_{0,i}^{*}\left\|\widehat{\bm{\mu}}\right\|,\sigma^{2}\big{)},\quad\widetilde{X}^{b}_{1,i}|Z_{1,i}^{*}\sim N\big{(}Z_{1,i}^{*}\left\|\widehat{\bm{\mu}}\right\|,\sigma^{2}\big{)},\quad\forall i\in[n],\quad\forall b\in[B].$
Similar to (24), we have $\textsf{TC}\left(\mathbf{X}_{0}^{b},\mathbf{X}_{1}^{b};\lambda,{\widehat{\bm{\mu}}},\widehat{\sigma}\right)=\textsf{TC}\big{(}\widetilde{}\mathbf{X}_{0}^{b},\widetilde{}\mathbf{X}_{1}^{b};\lambda,\left\|\widehat{\bm{\mu}}\right\|,\widehat{\sigma}\big{)}$ for any $\lambda>0$. Therefore, ^$\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}}$ given by ATC using $\big{(}\mathbf{X}_{0},\mathbf{X}_{1}\big{)}$ and $\left(\widehat{\bm{\mu}},\widehat{\sigma}\right)$ is exactly the same as the output of ATC using the projected samples (~$\big{(}\widetilde{}\mathbf{X}_{0},\widetilde{}\mathbf{X}_{1}\big{)}$ and $\left(\left\|\widehat{\bm{\mu}}\right\|,\widehat{\sigma}\right)$, i.e., ^$\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}}=\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,{\widehat{\bm{\mu}}},\widehat{\sigma}\right)=\textsf{ATC}\big{(}\widetilde{}\mathbf{X}_{0},\widetilde{}\mathbf{X}_{1};\Lambda,\left\|\widehat{\bm{\mu}}\right\|,\widehat{\sigma}\big{)}$. The problem is then reduced to the univariate case with samples $\big{\{}\widetilde{X}_{0,i},\widetilde{X}_{1,i}\big{\}}_{i=1}^{n}$ and unknown $\left(\widetilde{\mu},\sigma,\varepsilon\right)$.
We start by noting that $\delta_{n}=o\left(\frac{1}{\left(\left\|{\bm{\mu}}\right\|/\sigma\right)^{3}}\wedge\left(\frac{\left\|{\bm{\mu}}\right\|/\sigma}{\log n}\right)^{2}\right)$ implies that $\delta_{n}=o\left(\frac{1}{\left(\widetilde{\mu}/\sigma\right)^{3}}\wedge\left(\frac{\widetilde{\mu}/\sigma}{\log n}\right)^{2}\right)$ by 3, as
$\left|\frac{\widetilde{\mu}}{\mu}-1\right|=\frac{\left|{\widehat{\bm{\mu}}^{\top}{\bm{\mu}}}-\left\|\widehat{\bm{\mu}}\right\|\left\|{\bm{\mu}}\right\|\right|}{\left\|\widehat{\bm{\mu}}\right\|\left\|{\bm{\mu}}\right\|}\leq\frac{2\left\|\widehat{\bm{\mu}}-{\bm{\mu}}\right\|\left\|{\bm{\mu}}\right\|}{\left(1-\delta_{n}\right)\left\|{\bm{\mu}}\right\|^{2}}=O\left(\delta_{n}\right).$
We have the following lemma whose proof is deferred to Section D.9.
Lemma 11.
Under 3, 4 is satisfied for (~$\big{(}\widetilde{}\mathbf{X}_{0},\widetilde{}\mathbf{X}_{1}\big{)}$ with parameter $\left(\widetilde{\mu},\sigma,\varepsilon\right)$.
By Lemma 11 and Theorem 7, we immediately obtain that with probability at least $1-\zeta$,
(^$\ell\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{1}\min_{\lambda\in\Lambda}\left\{\mathcal{M}\left(\frac{\widetilde{\mu}}{\sigma},\varepsilon,\lambda\right)+\frac{\log\left(M\zeta^{-1}\right)}{n}\right\}.$
Recall that
$\mathcal{M}(s,\varepsilon,\lambda)=\Phi(-\sqrt{2}s)+\Phi\bigg{(}-s-\frac{\lambda}{2s}\bigg{)}\Phi\bigg{(}s-\frac{\lambda}{2s}\bigg{)}+2\varepsilon\Phi\bigg{(}-s+\frac{\lambda}{2s}\bigg{)}.$
It suffices to show that $\mathcal{M}\left(\widetilde{\mu}/\sigma,\varepsilon,\lambda\right)\lesssim\mathcal{M}\left(\mu/\sigma,\varepsilon,\lambda\right)$. By Lemma 5, we have
$\Phi\left(-\frac{\sqrt{2}\widetilde{\mu}}{\sigma}\right)\lesssim\frac{\widetilde{\mu}}{\mu}\exp\left(\frac{\left|\mu^{2}-\widetilde{\mu}^{2}\right|}{\sigma^{2}}\right)\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right).$
Moreover, using $\delta_{n}=O\left(\left(\log n\right)^{-1}\right)$ and $\mu^{2}/\sigma^{2}\lesssim\log n$, we have
$\exp\left(\frac{\mu^{2}-\widetilde{\mu}^{2}}{\sigma^{2}}\right)=\exp\left(\frac{\left(\mu-\widetilde{\mu}\right)\left(\mu+\widetilde{\mu}\right)}{\sigma^{2}}\right)\leq\exp\left(\frac{C\delta_{n}\mu^{2}}{\sigma^{2}}\right)\lesssim 1.$
Hence we obtain that $\Phi\left(-\sqrt{2}\widetilde{\mu}/\sigma\right)\lesssim\Phi\left(-\sqrt{2}\mu/\sigma\right)$. Then define
$\eta_{0}:={{\frac{\mu}{\sigma}-\frac{\lambda}{\mu/\sigma}}},\quad\Delta:=\left|-\frac{\widetilde{\mu}}{\sigma}+\frac{\lambda}{\widetilde{\mu}/\sigma}-\left(-\frac{\mu}{\sigma}+\frac{\lambda}{\mu/\sigma}\right)\right|$
Notice that 3 implies that $\Delta=O\left(\delta_{n}\left|\eta_{0}\right|\right)$. Hence by Lemma 6 we have
$\Phi\left(-\frac{\widetilde{\mu}}{\sigma}+\frac{\lambda}{\widetilde{\mu}/\sigma}\right)$ $\leq\Phi\left(-\eta_{0}\right)+\frac{C\delta_{n}\left|\eta_{0}\right|}{\sqrt{2\pi}}\exp\left(-\frac{\eta_{0}^{2}}{2}\right)\exp\left(C\delta_{n}\eta_{0}^{2}\right)$
$\lesssim\Phi\left(-\eta_{0}\right)+\frac{\delta_{n}\left|\eta_{0}\right|}{\sqrt{2\pi}}\exp\left(-\frac{\eta_{0}^{2}}{2}\right).$
where the second inequality holds due to $\delta_{n}=O\left(\left(\log n\right)^{-1}\right)$ and $\left|\eta_{0}\right|\lesssim\log n$. For $\eta_{0}>0$, we have
$\Phi\left(-\eta_{0}\right)\geq\frac{\sqrt{2/\pi}}{\eta_{0}+\sqrt{\eta_{0}^{2}+4}}\exp\left(-\frac{\eta_{0}^{2}}{2}\right)\gtrsim\frac{\delta_{n}\left|\eta_{0}\right|}{\sqrt{2\pi}}\exp\left(-\frac{\eta_{0}^{2}}{2}\right).$
For $\eta_{0}<0$, we have
$\frac{\delta_{n}\left|\eta_{0}\right|}{\sqrt{2\pi}}\exp\left(-\frac{\eta_{0}^{2}}{2}\right)<\frac{1}{2}<\Phi\left(-\eta_{0}\right).$
Thus we can conclude that
$\Phi\left(-\frac{\widetilde{\mu}}{\sigma}+\frac{\lambda}{\widetilde{\mu}/\sigma}\right)\lesssim\Phi\left(-\eta_{0}\right)=\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{\mu/\sigma}\right).$
Similarly, we could obtain that
$\Phi\left(\frac{\widetilde{\mu}}{\sigma}-\frac{\lambda}{\widetilde{\mu}/\sigma}\right)\lesssim\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{\mu/\sigma}\right),\quad\Phi\left(-\frac{\widetilde{\mu}}{\sigma}-\frac{\lambda}{\widetilde{\mu}/\sigma}\right)\lesssim\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{\mu/\sigma}\right).$
We thus conclude the proof by using the definition of $\mathcal{M}$.
Appendix C Other proofs
C.1 Proof of Corollary 1
We first show the following holds:
$\min_{\lambda\in\Lambda}\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\lambda\right)\lesssim\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right).$ (25)
Suppose that $\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)\in[\lambda_{j^{*}},\lambda_{j^{*}+1})$ for some $j^{*}\in[M]$. Denote
$\Delta:=\frac{\log n}{2M\mu/\sigma},\quad\eta_{0}:={{\frac{\mu}{\sigma}-\frac{\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)}{2\mu/\sigma}}}.$
If $\left|\eta_{0}\right|=o\left(\mu/\sigma\right)$ we have $\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)=\left(2\mu^{2}/\sigma^{2}\right)\left(1+o(1)\right)$ and hence
$\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)}{2\mu/\sigma}\right)\leq\exp\left(-\frac{2\mu^{2}}{\sigma^{2}}\left(1+o(1)\right)\right)=o\left(\Phi\left(-\sqrt{2}\mu/{\sigma}\right)\right).$
Similarly, we have
$\Phi\left(-\frac{\mu}{\sigma}-\frac{\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)}{2\mu/\sigma}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)}{2\mu/\sigma}\right)$
$\leq\Phi\left(-\frac{2\mu}{\sigma}\left(1+o(1)\right)\right)=o\left(\Phi\left(-\sqrt{2}\mu/{\sigma}\right)\right).$
Since $\Delta=o\left(\mu/\sigma\right)$, we can deduce that
$\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda_{j^{*}}}{2\mu/\sigma}\right)\leq\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)}{2\mu/\sigma}+\Delta\right)=o\left(\Phi\left(-\sqrt{2}\mu/{\sigma}\right)\right),$
$\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda_{j^{*}}}{2\mu/\sigma}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda_{j^{*}}}{2\mu/\sigma}\right)=o\left(\Phi\left(-\sqrt{2}\mu/{\sigma}\right)\right).$
Thus we have
$\min_{\lambda\in\Lambda}\left\{\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)+2\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right)\right\}$
$\leq\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda_{j^{*}}}{2\mu/\sigma}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda_{j^{*}}}{2\mu/\sigma}\right)+2\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda_{j^{*}}}{2\mu/\sigma}\right)=o\left(\Phi\left(-\sqrt{2}\mu/{\sigma}\right)\right).$
As a result of definition of $\mathcal{M}$, we obtain that
$\min_{\lambda\in\Lambda}\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\lambda\right)\leq\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right).$
Notice that the above argument also holds for any $\eta_{0}<0$, in which we have $\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)>2\mu^{2}/\sigma^{2}$, and $\Phi\left(-\sqrt{2}\mu/\sigma\right)$ is the dominant term. It suffices to consider $\eta_{0}\asymp\mu/\sigma$ ($\eta_{0}\leq\mu/\sigma$ by definition). Notice that in this regime, we have
$\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right)\geq\Phi\left(-\sqrt{2}\mu/\sigma\right)+\frac{1}{2}\Phi\left(-\frac{\mu}{\sigma}-\frac{\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)}{2\mu/\sigma}\right)+2\varepsilon\Phi\left(-\eta_{0}\right).$
On the other hand, by Lemma 6 we have
$\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda_{j^{*}+1}}{2\mu/\sigma}\right)\leq\Phi\left(-\eta_{0}+\Delta\right)\leq\Phi\left(-\eta_{0}\right)+\frac{\Delta}{\sqrt{2\pi}}\exp\left(-\frac{\left(\eta_{0}-\Delta\right)^{2}}{2}\right).$
Notice that $\Delta\exp\left(-{\left(\eta_{0}-\Delta\right)^{2}}/{2}\right)\asymp\Delta\exp\left(-{\eta_{0}^{2}}/{2}\right)$ and $\Phi\left(-\eta_{0}\right)\asymp{\eta_{0}}^{-1}\exp\left(-\frac{\eta_{0}^{2}}{2}\right)$, hence we have
$\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda_{j^{*}+1}}{2\mu/\sigma}\right)\lesssim\Phi\left(-\eta_{0}\right),$
provided that $\Delta=o\left(\left(\mu/\sigma\right)^{-1}\right)$, which is further implied by $M=\omega\left(\log n\right)$. Thus we have
$\min_{\lambda\in\Lambda}\left\{\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)+2\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right)\right\}$
$\leq\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda_{j^{*}+1}}{2\mu/\sigma}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda_{j^{*}+1}}{2\mu/\sigma}\right)+2\varepsilon\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda_{j^{*}+1}}{2\mu/\sigma}\right)$
$\lesssim\Phi\left(-\frac{\mu}{\sigma}-\frac{\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)}{2\mu/\sigma}\right)+\varepsilon\Phi\left(-\eta_{0}\right)\lesssim\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right).$
This concludes the proof for (25). It suffices to combine the (25), Theorem 4 and Theorem 2 to conclude that the additive term $\log(M\zeta^{-1})/n$ is ignorable provided that $\log\left(\zeta^{-1}\right)=O\left(\log n\right)$ and
$\textsf{SNR}\cdot\min\left\{\left(1+\alpha\right)^{2},2\right\}<\log n.$
C.2 Proof of Proposition 1
Fix $i\in[n]$. We have $\{X_{0,i}>0\}\subseteq\{\check{Z}^{\textsf{ITL}}_{0,i}=1\}\subseteq\{X_{0,i}\geq 0\}$. Then,
$\mathbb{P}(\check{Z}^{\textsf{ITL}}_{0,i}\neq Z^{*}_{0,i}\mid Z^{*}_{0,i}=1)$
$=\mathbb{P}\left(X_{0,i}<0\mid Z^{*}_{0,i}=1\right)=\mathbb{P}\bigg{(}\frac{X_{0,i}-\mu}{\sigma}<-\frac{\mu}{\sigma}\bigg{|}Z^{*}_{0,i}=1\bigg{)}=\Phi\left(-\frac{\mu}{\sigma}\right).$
By symmetry, we can obtain the desired result for $\mathbb{P}(\check{Z}^{\textsf{ITL}}_{0,i}\neq Z^{*}_{0,i})$.
It remains to study $\check{Z}^{\textsf{DP}}_{0,i}$. From $\{X_{0,i}+X_{1,i}>0\}\subseteq\{\check{Z}^{\textsf{DP}}_{0,i}=1\}\subseteq\{X_{0,i}+X_{1,i}\geq 0\}$ we obtain that
$\mathbb{P}_{1,1}\left(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i}\right)$ $=\mathbb{P}_{1,1}\left(X_{0,i}+X_{1,i}<0\right)$
$=\mathbb{P}_{1,1}\bigg{(}\frac{X_{0,i}+X_{1,i}-2\mu}{\sqrt{2}\sigma}<-\sqrt{2}\cdot\frac{\mu}{\sigma}\bigg{)}=\Phi\bigg{(}-\frac{\sqrt{2}\mu}{\sigma}\bigg{)}.$ (26)
On the other hand,
$\mathbb{P}_{1,-1}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i})$ $=\mathbb{P}_{1,-1}\left(X_{0,i}+X_{1,i}<0\right)=\mathbb{P}_{1,-1}\left(\frac{X_{0,i}+X_{1,i}}{\sqrt{2}\sigma}<0\right)=\frac{1}{2}.$ (27)
By symmetry, the right-hand sides of (26) and (27) can upper bound $\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}=Z^{*}_{1,i})$ and $\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}\neq Z^{*}_{1,i})$, respectively. Therefore,
$\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i})$
$=\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}=Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}=Z^{*}_{1,i})+\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i}|Z^{*}_{0,i}\neq Z^{*}_{1,i})\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i})$
$=\Phi\bigg{(}-\frac{\sqrt{2}\mu}{\sigma}\bigg{)}\cdot[1-\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i})]+\frac{1}{2}\cdot\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i})$
$=\Phi\bigg{(}-\frac{\sqrt{2}\mu}{\sigma}\bigg{)}+\bigg{[}\frac{1}{2}-\Phi\bigg{(}-\frac{\sqrt{2}\mu}{\sigma}\bigg{)}\bigg{]}\mathbb{P}(Z^{*}_{0,i}\neq Z^{*}_{1,i}).$
We have
$0\leq\mathbb{P}(\check{Z}^{\textsf{DP}}_{0,i}\neq Z^{*}_{0,i})-\Phi\bigg{(}-\frac{\sqrt{2}\mu}{\sigma}\bigg{)}\leq\frac{\varepsilon}{2}.$
Appendix D Proofs of Lemmas
D.1 Proof of Lemma 1
Without loss of generality, we assume $\sigma=1$. Notice that for any $i\in[n]$, we have
$\mathbb{I}\left(\check{Z}^{\textsf{ITL}}_{0,i}\neq\check{Z}^{\textsf{ITL}}_{1,i}\right)$ $=\mathbb{I}\left(Z^{*}_{0,i}\neq Z^{*}_{1,i},\check{Z}^{\textsf{ITL}}_{0,i}=Z^{*}_{0,i},\check{Z}^{\textsf{ITL}}_{1,i}=Z^{*}_{1,i}\right)$
$+\mathbb{I}\left(Z^{*}_{0,i}\neq Z^{*}_{1,i},\check{Z}^{\textsf{ITL}}_{0,i}\neq Z^{*}_{0,i},\check{Z}^{\textsf{ITL}}_{1,i}\neq Z^{*}_{1,i}\right)$
$+\mathbb{I}\left(Z^{*}_{0,i}=Z^{*}_{1,i},\check{Z}^{\textsf{ITL}}_{0,i}=Z^{*}_{0,i},\check{Z}^{\textsf{ITL}}_{1,i}\neq Z^{*}_{1,i}\right)$
$+\mathbb{I}\left(Z^{*}_{0,i}=Z^{*}_{1,i},\check{Z}^{\textsf{ITL}}_{0,i}\neq Z^{*}_{0,i},\check{Z}^{\textsf{ITL}}_{1,i}=Z^{*}_{1,i}\right).$
Averaging over $i\in[n]$, taking expectations on both sides and using Proposition 1, we obtain that
$\mathbb{E}\left(\widehat{\varepsilon}\right)$ $=\mathbb{P}\left(\check{Z}^{\textsf{ITL}}_{0,i}\neq\check{Z}^{\textsf{ITL}}_{1,i}\right)$
$=\varepsilon\cdot\left[\left[1-\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}+\left[\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}\right]+2\left(1-\varepsilon\right)\Phi\left(-\frac{\mu}{\sigma}\right)\left[1-\Phi\left(-\frac{\mu}{\sigma}\right)\right].$
We thus get
$\left|\mathbb{E}\left(\widehat{\varepsilon}\right)-\varepsilon\right|$ $=2\varepsilon\cdot\left[1+\Phi\left(-\frac{\mu}{\sigma}\right)\right]\Phi\left(-\frac{\mu}{\sigma}\right)+2\left(1-\varepsilon\right)\Phi\left(-\frac{\mu}{\sigma}\right)\left[1-\Phi\left(-\frac{\mu}{\sigma}\right)\right]$
$=2\left[1-\left(1-2\varepsilon\right)\Phi\left(-\frac{\mu}{\sigma}\right)\right]\Phi\left(-\frac{\mu}{\sigma}\right)\geq\Phi\left(-\frac{\mu}{\sigma}\right).$
Moreover, since $\mathbb{P}\left(\check{Z}^{\textsf{ITL}}_{0,i}\neq\check{Z}^{\textsf{ITL}}_{1,i}\right)<1/2$ and $\Phi\left(-\mu/\sigma\right)<1/4$ for $\mu/\sigma\rightarrow\infty$, we have
$\text{Var}\left(\widehat{\varepsilon}\right)$ $=\frac{1}{n^{2}}\sum_{i=1}^{n}\mathbb{P}\left(\check{Z}^{\textsf{ITL}}_{0,i}\neq\check{Z}^{\textsf{ITL}}_{1,i}\right)\left(1-\mathbb{P}\left(\check{Z}^{\textsf{ITL}}_{0,i}\neq\check{Z}^{\textsf{ITL}}_{1,i}\right)\right)$
$\geq\frac{1}{2n}\mathbb{P}\left(\check{Z}^{\textsf{ITL}}_{0,i}\neq\check{Z}^{\textsf{ITL}}_{1,i}\right)\geq\frac{1}{4n}\left[\varepsilon+\Phi\left(-\frac{\mu}{\sigma}\right)\right].$
Therefore, we have
$\mathbb{E}\left(\widehat{\varepsilon}-\varepsilon\right)^{2}$ $=\text{Var}\left(\widehat{\varepsilon}\right)+\left(\mathbb{E}\left(\widehat{\varepsilon}\right)-\varepsilon\right)^{2}$
$\geq\frac{1}{4n}\left[\varepsilon+\Phi\left(-\frac{\mu}{\sigma}\right)\right]+\left[\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}.$
We thus arrive at
$\mathbb{E}\left(\frac{\widehat{\varepsilon}}{\varepsilon}-1\right)^{2}$
$\geq\frac{1/\varepsilon}{4n}+\frac{1/\varepsilon^{2}}{4n}\cdot\Phi\left(-\frac{\mu}{\sigma}\right)+\left[\frac{1}{\varepsilon}\cdot\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}$
$\gtrsim\frac{1}{n^{1-\beta}}+\frac{1}{\mu/\sigma}\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}+\left(2\beta-1\right)\log n\right)+\left[\frac{1}{\mu/\sigma}\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}+\beta\log n\right)\right]^{2}.$
The desired result follows directly.
D.2 Proof of Lemma 2
First note that for any $\lambda\in\Lambda$ and $\lambda^{\prime}<\lambda$ we have
${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\widehat{}\mathbf{Z}_{0}^{\lambda^{\prime}}\right)-\widehat{\psi}\left(\lambda^{\prime}\right)$
$\leq{\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)+{\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda^{\prime}},\bar{}\mathbf{Z}_{0}^{\infty}\right)-\widehat{\psi}\left(\lambda^{\prime}\right)$
$\leq{\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)+{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)+{\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda^{\prime}},\bar{}\mathbf{Z}_{0}^{\lambda^{\prime}}\right)+{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda^{\prime}},\bar{}\mathbf{Z}_{0}^{\infty}\right)-\widehat{\psi}\left(\lambda^{\prime}\right)$
$\leq\phi\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)+\xi_{\widehat{\psi}}+\phi\left(\lambda^{\prime}\right)+{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda^{\prime}},\bar{}\mathbf{Z}_{0}^{\infty}\right)-\widehat{\psi}\left(\lambda^{\prime}\right).$
Taking maximum over $\lambda^{\prime}\in[0,\lambda)\bigcap\Lambda$ leads to
$0\leq\widehat{\phi}\left(\lambda\right)\leq 2\phi\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)+2\xi_{\widehat{\psi}},\quad\forall\lambda\in\Lambda.$ (28)
For any $\lambda\in[0,\widehat{\lambda})\bigcap\Lambda$, we have
${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ $\leq{\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\widehat{}\mathbf{Z}_{0}^{\lambda}\right)+{\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)$
$\leq\left[\widehat{\phi}\left(\widehat{\lambda}\right)+\widehat{\psi}\left(\lambda\right)\right]+\left[\phi\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)+\xi_{\widehat{\psi}}\right]$ (definition of $(\text{definition of }\widehat{\phi})$
$\leq\left[\widehat{\phi}\left(\widehat{\lambda}\right)+\widehat{\psi}\left(\widehat{\lambda}\right)\right]+\left[\phi\left(\lambda\right)+2\widehat{\psi}\left(\lambda\right)+\xi_{\widehat{\psi}}\right]$
$\leq\left[\widehat{\phi}\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right]+\left[\phi\left(\lambda\right)+2\widehat{\psi}\left(\lambda\right)+\xi_{\widehat{\psi}}\right]$ (definition of $(\text{definition of }\widehat{\lambda})$
[…]
For any $\lambda\in[\widehat{\lambda},\infty)\bigcap\Lambda$, we have
${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ $\leq{\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\widehat{\lambda}}\right)+{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)$
$\leq\phi\left(\widehat{\lambda}\right)+\widehat{\psi}\left(\widehat{\lambda}\right)+\xi_{\widehat{\psi}}$ (definition of $(\text{definition of }\widehat{\phi}\text{~{}and~{}}\xi_{\widehat{\psi}})$
$\leq\phi\left(\lambda\right)+\left[\widehat{\phi}\left(\widehat{\lambda}\right)+\widehat{\psi}\left(\widehat{\lambda}\right)\right]+\xi_{\widehat{\psi}}$
$\leq\phi\left(\lambda\right)+\left[\widehat{\phi}\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right]+\xi_{\widehat{\psi}}$ (definition of $(\text{definition of }\widehat{\lambda})$
[…]
D.3 Proof of Lemma 3
For each $i\in[n]$ we have
$\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq\bar{Z}^{\infty}_{0,i}\right)$ $\leq\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq\widetilde{Z}^{\lambda}_{0,i}\right)+\mathbb{P}\left(\widetilde{Z}^{\lambda}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}\right)+\mathbb{P}\left(\widetilde{Z}^{\infty}_{0,i}\neq\bar{Z}^{\infty}_{0,i}\right)$
[…]
$\leq\mathbb{P}\left(\widetilde{Z}^{\lambda}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}\right)+\gamma\left[\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq\bar{Z}_{0,i}^{\infty}\right)+2\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq Z_{0,i}^{*}\right)\right].$
Rearranging terms leads to
$\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq\bar{Z}^{\infty}_{0,i}\right)\leq\frac{1}{1-\gamma}\mathbb{P}\left(\widetilde{Z}^{\lambda}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}\right)+\frac{2\gamma}{1-\gamma}\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq Z_{0,i}^{*}\right).$ (29)
On the other hand, we have
$\mathbb{P}\left(\widetilde{Z}^{\lambda}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}\right)$ $\leq\mathbb{P}\left(\widetilde{Z}^{\lambda}_{0,i}\neq\bar{Z}^{\lambda}_{0,i}\right)+\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq\bar{Z}^{\infty}_{0,i}\right)+\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}\right)$
[…]
$\leq\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq\bar{Z}^{\infty}_{0,i}\right)+\gamma\left[\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq\bar{Z}_{0,i}^{\infty}\right)+2\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq Z_{0,i}^{*}\right)\right]$
$=\left(1+\gamma\right)\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq\bar{Z}^{\infty}_{0,i}\right)+2\gamma\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq Z_{0,i}^{*}\right).$ (30)
We need the following data-driven bound for two sample mean of Bernoulli r.v’s, whose proof is given in Section D.10.
Lemma 12.
Let $\left\{X_{0,i}\right\}_{i=1}^{n}$ be i.i.d. Ber$\left(p\right)$, $\left\{X_{1,i}\right\}_{i=1}^{n}$ be i.i.d. Ber$\left(q\right)$, and $\left\{X_{0,i}\right\}_{i=1}^{n}$ are independent of $\left\{X_{1,i}\right\}_{i=1}^{n}$. Suppose that $p\leq\vartheta_{1}q+\vartheta_{2}$ for some $\vartheta_{1},\vartheta_{2}>0$, then for any $\vartheta\in(0,1],\delta\in(0,1)$ we have
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{0,i}\leq\left(1+\vartheta\right)\vartheta_{1}\left(\frac{1}{n}\sum_{i=1}^{n}X_{1,i}\right)+\frac{7\left(1+2\vartheta_{1}\right)\log\left(2/\delta\right)}{2\vartheta n}+2\vartheta_{2}\right)\geq 1-\delta.$
Invoking Lemma 12 for any $\vartheta,\zeta\in(0,1)$, we obtain that with probability at least $1-\zeta$,
${\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)\leq\frac{1+\vartheta}{1-\gamma}{\mathfrak{D}}\left(\widetilde{}\mathbf{Z}^{\lambda}_{0},\widetilde{}\mathbf{Z}^{\infty}_{0}\right)+\left(\frac{3-\gamma}{1-\gamma}\right)\frac{7\log\left(2/\zeta\right)}{2\vartheta n}+\frac{4\gamma}{1-\gamma}\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right),$ (31)
and with probability at least $1-\zeta$,
${\mathfrak{D}}\left(\widetilde{}\mathbf{Z}^{\lambda}_{0},\widetilde{}\mathbf{Z}^{\infty}_{0}\right)\leq\left(1+\vartheta\right)\left(1+\gamma\right){\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)+\left(3+2\gamma\right)\frac{7\log\left(2/\zeta\right)}{2\vartheta n}+4\gamma\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right).$ (32)
Let $\bar{Q}_{\zeta}^{\lambda}:=\left(1-\zeta\right)$ population quantile of ${\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda,1}_{0},\bar{}\mathbf{Z}^{\infty,1}_{0}\right)$ conditional on $(\widehat{\mu},\widehat{\sigma})$.
Define the quantile function as $Q_{\zeta}(F):=\inf\left\{z\in\mathbb{R}:F(z)\geq\zeta\right\}$ for any $\zeta\in(0,1)$ and any c.d.f. $F$ on $\mathbb{R}$. The following lemma is needed to control the deviation of $\bar{Q}_{\zeta}^{\lambda}$ to $\widehat{Q}_{\zeta}^{\lambda}$.
Lemma 13.
Suppose $X,X_{1},X_{2},\cdots,X_{N}\overset{i.i.d.}{\sim}F$ where $F\left(\cdot\right)$ is the c.d.f., and denote the empirical c.d.f. of $\left\{X_{i}\right\}_{i=1}^{N}$ as $\widehat{F}_{N}\left(\cdot\right)$. For any $\zeta,\eta\in(0,1)$ we have with probability at least $1-\eta$,
$Q_{1-\zeta}\left(F\right)\leq Q_{1-{\zeta}/{2}}(\widehat{F}_{N})\leq Q_{1-{\zeta}/{3}}\left(F\right),$
provided that $N\gtrsim\zeta^{-2}\log\left(1/\eta\right)$.
By Lemma 13 and a union bound, we can conclude that with probability at least $1-\zeta$,
$\bar{Q}_{\zeta}^{\lambda}\leq\widehat{Q}_{\zeta/2}^{\lambda}\leq\bar{Q}_{\zeta/3}^{\lambda},\quad\forall\lambda\in\Lambda$ (33)
provided that $B\geq C\zeta^{-2}\log\left(\left|\Lambda\right|/\zeta\right)$ for some large constant $C>0$. Hence we
By definition of $\bar{Q}^{\lambda}_{\zeta}$ we have
$\mathbb{P}\left({\mathfrak{D}}\left(\widetilde{}\mathbf{Z}^{\lambda}_{0},\widetilde{}\mathbf{Z}^{\infty}_{0}\right)<\bar{Q}_{\zeta}^{\lambda}\right)\leq 1-\zeta,\quad\mathbb{P}\left({\mathfrak{D}}\left(\widetilde{}\mathbf{Z}^{\lambda}_{0},\widetilde{}\mathbf{Z}^{\infty}_{0}\right)\leq\bar{Q}_{\zeta}^{\lambda}\right)\geq 1-\zeta.$ (34)
Combined with (31) and (33), we obtain that with probability at least $1-3\zeta$,
${\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)\leq\frac{1+\vartheta}{1-\gamma}\widehat{Q}_{\zeta/2}^{\lambda}+\left(\frac{3-\gamma}{1-\gamma}\right)\frac{7\log\left(4/\zeta\right)}{2\vartheta n}+\frac{4\gamma}{1-\gamma}\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right).$
By choosing $\vartheta=0.005$ and noting $\gamma<0.04$, we thus obtain that with probability at least $1-\zeta$,
${\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)\leq\widehat{\psi}\left(\lambda\right)+2200\left[\frac{\log\left(1/\zeta\right)}{n}+\gamma\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right)\right].$ (35)
By definition of $\xi_{\widehat{\psi}}$ we conclude the proof of the first claim.
It suffices to prove the upper bound for $\widehat{\psi}\left(\lambda\right)$. By (32) with $\vartheta=1/2$, we obtain that with probability at least $61-\zeta/$ 6,
${\mathfrak{D}}\left(\widetilde{}\mathbf{Z}^{\lambda}_{0},\widetilde{}\mathbf{Z}^{\infty}_{0}\right)\leq 3{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)+\frac{35\log\left(12/\zeta\right)}{n}+4\gamma\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right).$
By Lemma 4 with $\vartheta=1$, we obtain that with probability at least $61-\zeta/$ 6,
${\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)\leq 2\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)+\frac{7\log\left(6/\zeta\right)}{6n}$
By a union bound, we conclude that with probability at least $31-\zeta/$ 3,
${\mathfrak{D}}\left(\widetilde{}\mathbf{Z}^{\lambda}_{0},\widetilde{}\mathbf{Z}^{\infty}_{0}\right)\leq 6\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)+\frac{40\log\left(12/\zeta\right)}{n}+4\gamma\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right).$
Together with (33) and (34), we can conclude that with probability at least $31-10\zeta/$ 3,
$\widehat{\psi}\left(\lambda\right)=1.01\widehat{Q}_{\zeta/2}^{\lambda}\leq 1.01\bar{Q}_{\zeta/3}^{\lambda}\leq 45\left[\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\lambda}_{0},\bar{}\mathbf{Z}^{\infty}_{0}\right)+\frac{\log\left(1/\zeta\right)}{n}+\gamma\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right)\right].$ (36)
The proof is then completed by adjusting the constant in front of $\zeta$ and taking a union bound over (35) and (36).
D.4 Proof of Lemma 4
By Bernstein’s inequality, see, e.g., Lemma 3.1 in [27], we have with probability at least $1-\delta$ that
$\frac{1}{n}\sum_{i=1}^{n}X_{0,i}-p\leq\sqrt{\frac{2p\log\left(1/\delta\right)}{n}}+\frac{2\log\left(1/\delta\right)}{3n}=\inf_{\vartheta>0}\left\{\vartheta p+\frac{\log\left(1/\delta\right)}{2\vartheta n}\right\}+\frac{2\log\left(1/\delta\right)}{3n}.$
This implies for any $\vartheta\in(0,1]$, we have with probability at least $1-\delta$ that
$\frac{1}{n}\sum_{i=1}^{n}X_{0,i}\leq\left(1+\vartheta\right)p+\left(\frac{1}{2\vartheta}+\frac{2}{3}\right)\frac{\log\left(1/\delta\right)}{n}\leq\left(1+\vartheta\right)p+\frac{7\log\left(1/\delta\right)}{6\vartheta n}.$
The lower bound can be obtained similarly and hence omitted.
D.5 Proof of Lemma 7
Denote $K^{+}=\frac{\lambda\sigma^{2}}{2\mu}\left(1+\delta_{n}^{\prime}\right)$ and $K^{-}=\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}^{\prime}\right)$. By symmetry, it suffices for us to consider the case when $(Z^{*}_{0,i},Z^{*}_{1,i})=(1,1)$. First, note that the transfer clustering procedure would yield $\widehat{Z}^{\lambda}_{0,i}=-1$ if
$X_{0,i}<X_{1,i},\quad X_{0,i}<-\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}},\quad X_{1,i}>\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}}$ (37)
or
$X_{0,i}+X_{1,i}<0,\quad X_{0,i}<\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}},\quad X_{1,i}<\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}}.$ (38)
We have
[…]
$=\mathbb{P}_{1,1}\left(X_{0,i}<-\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}},X_{1,i}>\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}},{\cal E}_{\sf par}\right)$
$\geq\mathbb{P}_{1,1}\left(X_{0,i}<-K^{+}\right)\mathbb{P}_{1,1}\left(X_{1,i}>K^{+}\right)$
$=\Phi\left(\frac{-\mu-K^{+}}{\sigma}\right)\left(1-\Phi\left(\frac{-\mu+K^{+}}{\sigma}\right)\right).$
It remains to lower bound the following term
[…]
$\geq\mathbb{P}_{1,1}\left(X_{0,i}+X_{1,i}<0,X_{0,i}<\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}^{\prime}\right),X_{1,i}<\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}^{\prime}\right)\right)$
$=1-\mathbb{P}_{1,1}\left(\left\{X_{0,i}+X_{1,i}\geq 0\right\}\bigcup\left\{X_{0,i}\geq K^{-}\right\}\bigcup\left\{X_{1,i}\geq K^{-}\right\}\right)$
$=1-\mathbb{P}_{1,1}\left(X_{0,i}+X_{1,i}\geq 0\right)-\mathbb{P}_{1,1}\left(X_{0,i}\geq K^{-}\right)-\mathbb{P}_{1,1}\left(X_{1,i}\geq K^{-}\right)$
$+\mathbb{P}_{1,1}\left(\left\{X_{0,i}+X_{1,i}\geq 0\right\}\bigcap\left\{X_{0,i}\geq K^{-}\right\}\right)+\mathbb{P}_{1,1}\left(\left\{X_{0,i}+X_{1,i}\geq 0\right\}\bigcap\left\{X_{1,i}\geq K^{-}\right\}\right)$
$=\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)-\mathbb{P}_{1,1}\left(\left\{X_{0,i}+X_{1,i}<0\right\}\bigcap\left\{X_{0,i}\geq K^{-}\right\}\right)$
$-\mathbb{P}_{1,1}\left(\left\{X_{0,i}+X_{1,i}<0\right\}\bigcap\left\{X_{1,i}\geq K^{-}\right\}\right).$
Notice that
$\mathbb{P}$ ($\left(\left\{X_{0,i}+X_{1,i}<0\right\}\bigcap\left\{X_{0,i}\geq K^{-}\right\}\right)$
[…]
$=\frac{1}{\sqrt{2}}\frac{\sqrt{2/\pi}\exp\left(-\mu^{2}/\sigma^{2}\right)}{\sqrt{2}\mu/\sigma+\sqrt{2\mu^{2}/\sigma^{2}+4}}\cdot\frac{\sqrt{2}\mu/\sigma+\sqrt{2\mu^{2}/\sigma^{2}+4}}{\left(K^{-}+\mu\right)/\sigma+\sqrt{\left(K^{-}+\mu\right)^{2}/\sigma^{2}+8/\pi}}\Phi\left(-\frac{\sqrt{2}K^{-}}{\sigma}\right)$
[…]
Another trivial bound can be obtained by
[…]
$\geq\mathbb{P}\left(X_{0,i}+X_{1,i}<0,X_{0,i}<\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}^{\prime}\right),X_{1,i}<\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}^{\prime}\right)\right)$
$\geq\mathbb{P}\left(X_{0,i}+X_{1,i}<0,X_{0,i}<0,X_{1,i}<0\right)=\left[\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}$
[…]
$\mathbb{P}\left(\widehat{Z}^{\lambda}_{0,i}=-1\bigg{|}(Z^{*}_{0,i},Z^{*}_{1,i})=(1,1)\right)$
$\geq\Phi\left(\frac{-\mu-K^{+}}{\sigma}\right)\Phi\left(\frac{\mu-K^{+}}{\sigma}\right)+\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)$
$+\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)\left(1-2\Phi\left(-\frac{\sqrt{2}K^{-}}{\sigma}\right)\right)\vee\left[\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}.$
By symmetry, a lower bound for $\mathbb{P}\left(\widehat{Z}^{\lambda}_{0,i}=1\bigg{|}(Z^{*}_{0,i},Z^{*}_{1,i})=(-1,-1)\right)$ can be derived similarly and hence omitted.
D.6 Proof of Lemma 8
Fix $i\in[n]$ and we can write
$\begin{cases}X_{0,i}=\mu Z^{*}_{0,i}+\sigma V_{0,i}\\ X_{1,i}=\mu Z^{*}_{1,i}+\sigma V_{1,i}\\ \bar{X}_{1,i}=\mu Z^{*}_{0,i}+\sigma V_{1,i}\end{cases},$ (39)
[…]
Bound for $\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq\bar{Z}^{\lambda}_{0,i}\right)$
By definition and (39), we have
$\bar{Z}^{\infty}_{0,i}$ $=Z^{*}_{0,i}\cdot\text{sign}\left(1+\frac{\sigma}{\sqrt{2}\mu}\frac{\left(V_{0,i}+V_{1,i}\right)Z^{*}_{0,i}}{\sqrt{2}}\right),$
$\left(\bar{Z}^{\lambda}_{0,i},\bar{Z}^{\lambda}_{1,i}\right)$ $\in\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\left(\frac{\mu}{\sigma}\right)^{2}Z^{*}_{0,i}\left(u+v\right)-\frac{\mu}{\sigma}\left(uV_{0,i}+vV_{1,i}\right)+\lambda\mathbb{I}\left(u\neq v\right)\right\}.$
Hence we have $\bar{Z}^{\infty}_{0,i}=1$ if
$Z^{*}_{0,i}=1,\quad V_{0,i}+V_{1,i}>-\frac{2\mu}{\sigma},$
or
$Z^{*}_{0,i}=-1,\quad V_{0,i}+V_{1,i}>\frac{2\mu}{\sigma}.$
On the other hand, we have $\bar{Z}^{\lambda}_{0,i}=-1$ if
$V_{0,i}<V_{1,i},\quad V_{0,i}<-\frac{\mu}{\sigma}Z^{*}_{0,i}-\frac{\lambda\sigma}{2\mu},\quad V_{1,i}>-\frac{\mu}{\sigma}Z^{*}_{0,i}+\frac{\lambda\sigma}{2\mu},$
or
$V_{0,i}+V_{1,i}<-\frac{2\mu}{\sigma}Z^{*}_{0,i},\quad V_{0,i}<-\frac{\mu}{\sigma}Z^{*}_{0,i}+\frac{\lambda\sigma}{2\mu},\quad V_{1,i}<-\frac{\mu}{\sigma}Z^{*}_{0,i}+\frac{\lambda\sigma}{2\mu}.$
Hence we have $\left(\bar{Z}^{\infty}_{0,i},\bar{Z}^{\lambda}_{0,i}\right)=\left(1,-1\right)$ only if
$Z^{*}_{0,i}=1,\quad V_{0,i}+V_{1,i}>-\frac{2\mu}{\sigma},\quad V_{0,i}<-\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu},\quad V_{1,i}>-\frac{\mu}{\sigma}+\frac{\lambda\sigma}{2\mu},$
or
$Z^{*}_{0,i}=-1,\quad V_{0,i}+V_{1,i}>\frac{2\mu}{\sigma},\quad V_{0,i}<\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu},\quad V_{1,i}>\frac{\mu}{\sigma}+\frac{\lambda\sigma}{2\mu}.$
We can conclude that
$\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}=1,\bar{Z}^{\lambda}_{0,i}=-1\right)$ $\leq 2\min\left\{\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right),\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)\right\}.$
By symmetry the bound for $\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}=-1,\bar{Z}^{\lambda}_{0,i}=1\right)$ is the same. Hence, we have obtained the first claim in Lemma 8.
Bound for $\mathbb{P}\left(\check{Z}_{0,i}^{\lambda}\neq\bar{Z}_{0,i}^{\lambda}\right)$
By definition and (39), we have
$\left(\check{Z}^{\lambda}_{0,i},\check{Z}^{\lambda}_{1,i}\right)$ $\in\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\left(\frac{\mu}{\sigma}\right)^{2}\left(Z^{*}_{0,i}u+Z^{*}_{1,i}v\right)-\frac{\mu}{\sigma}\left(uV_{0,i}+vV_{1,i}\right)+\lambda\mathbb{I}\left(u\neq v\right)\right\},$
$\left(\bar{Z}^{\lambda}_{0,i},\bar{Z}^{\lambda}_{1,i}\right)$ $\in\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\left(\frac{\mu}{\sigma}\right)^{2}Z^{*}_{0,i}\left(u+v\right)-\frac{\mu}{\sigma}\left(uV_{0,i}+vV_{1,i}\right)+\lambda\mathbb{I}\left(u\neq v\right)\right\}.$
Hence we have $\check{Z}^{\lambda}_{0,i}=1$ if
$V_{0,i}>V_{1,i}+\frac{\mu}{\sigma}\left(Z^{*}_{1,i}-Z^{*}_{0,i}\right),\quad V_{0,i}>-\frac{\mu}{\sigma}Z^{*}_{0,i}+\frac{\lambda\sigma}{2\mu},\quad V_{1,i}<-\frac{\mu}{\sigma}Z^{*}_{1,i}-\frac{\lambda\sigma}{2\mu},$
or
$V_{0,i}+V_{1,i}>-\frac{\mu}{\sigma}\left(Z^{*}_{0,i}+Z^{*}_{1,i}\right),\quad V_{0,i}>-\frac{\mu}{\sigma}Z^{*}_{0,i}-\frac{\lambda\sigma}{2\mu},\quad V_{1,i}>-\frac{\mu}{\sigma}Z^{*}_{1,i}-\frac{\lambda\sigma}{2\mu}.$
On the other hand, we have $\bar{Z}^{\lambda}_{0,i}=-1$ if
$V_{0,i}<V_{1,i},\quad V_{0,i}<-\frac{\mu}{\sigma}Z^{*}_{0,i}-\frac{\lambda\sigma}{2\mu},\quad V_{1,i}>-\frac{\mu}{\sigma}Z^{*}_{0,i}+\frac{\lambda\sigma}{2\mu},$
or
$V_{0,i}+V_{1,i}<-\frac{2\mu}{\sigma}Z^{*}_{0,i},\quad V_{0,i}<-\frac{\mu}{\sigma}Z^{*}_{0,i}+\frac{\lambda\sigma}{2\mu},\quad V_{1,i}<-\frac{\mu}{\sigma}Z^{*}_{0,i}+\frac{\lambda\sigma}{2\mu}.$
By definition, we have $\left(\check{Z}^{\lambda}_{0,i},\bar{Z}^{\lambda}_{0,i}\right)=\left(1,-1\right)$ only if $Z^{*}_{0,i}\neq Z^{*}_{1,i}$. When $\left(Z^{*}_{0,i},Z^{*}_{1,i}\right)=(1,-1)$, then (i) the first condition of $\check{Z}^{\lambda}_{0,i}=1$ conflicts with the conditions of $\bar{Z}^{\lambda}_{0,i}=-1$ due to the condition on $V_{0,i}$, (ii) the second condition of $\check{Z}^{\lambda}_{0,i}=1$ conflicts with the first condition of $\bar{Z}^{\lambda}_{0,i}=-1$ due to the condition on $V_{0,i}$ and conflicts with the second condition of $\bar{Z}^{\lambda}_{0,i}=-1$ due to the condition on $V_{0,i}+V_{1,i}$. Hence we have $\left(\check{Z}^{\lambda}_{0,i},\bar{Z}^{\lambda}_{0,i}\right)=\left(1,-1\right)$ only if $\left(Z^{*}_{0,i},Z^{*}_{1,i}\right)=(-1,1)$ and
$0<V_{0,i}+V_{1,i}<\frac{2\mu}{\sigma},\quad\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}<V_{0,i}<\frac{\mu}{\sigma}+\frac{\lambda\sigma}{2\mu},\quad-\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}<V_{1,i}<\frac{\mu}{\sigma}+\frac{\lambda\sigma}{2\mu},$
Hence we arrive at
$\mathbb{P}\left(\check{Z}^{\lambda}_{0,i}=1,\bar{Z}^{\lambda}_{0,i}=-1\right)$
$\leq\frac{\varepsilon}{2}\cdot\mathbb{P}\left(\check{Z}^{\lambda}_{0,i}=1,\bar{Z}^{\lambda}_{0,i}=-1\mid\left(Z^{*}_{0,i},Z^{*}_{1,i}\right)=(-1,1)\right)$
$\leq\frac{\varepsilon}{2}\cdot\mathbb{P}\left(\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}<V_{0,i}<\frac{\mu}{\sigma}+\frac{\lambda\sigma}{2\mu},\quad-\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}<V_{1,i}<-\frac{\mu}{\sigma}+\frac{\lambda\sigma}{2\mu}\right)$
$\leq\frac{\varepsilon}{2}\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda\sigma}{2\mu}\right)\left[1-\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)\right].$
By symmetry, we have the same bound for $\mathbb{P}\left(\check{Z}^{\lambda}_{0,i}=-1,\bar{Z}^{\lambda}_{0,i}=1\right)$ and the proof for the second claim of Lemma 8 is completed.
D.7 Proof of Lemma 9
For $\forall i\in[n]$, by definition we have
$\left(\check{Z}^{\lambda}_{0,i},\check{Z}^{\lambda}_{1,i}\right)$ $\in\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\mu\left(uX_{0,i}+vX_{1,i}\right)+\lambda\sigma^{2}\mathbb{I}\left(u\neq v\right)\right\}.$
Hence we have $\check{Z}^{\lambda}_{0,i}=1$ if
$X_{0,i}>X_{1,i},\quad X_{0,i}>\frac{\lambda\sigma^{2}}{2\mu},\quad X_{1,i}<-\frac{\lambda\sigma^{2}}{2\mu},$
or
$X_{0,i}+X_{1,i}>0,\quad X_{0,i}>-\frac{\lambda\sigma^{2}}{2\mu},\quad X_{1,i}>-\frac{\lambda\sigma^{2}}{2\mu}.$
On the other hand, we have $\widehat{Z}^{\lambda}_{0,i}=-1$ if
$X_{0,i}<X_{1,i},\quad X_{0,i}<-\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}},\quad X_{1,i}>\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}},$
or
$X_{0,i}+X_{1,i}<0,\quad X_{0,i}<\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}},\quad X_{1,i}<\frac{\lambda\widehat{\sigma}^{2}}{2\widehat{\mu}}.$
Hence under Assumption 4 we have $\left(\check{Z}^{\lambda}_{0,i},\widehat{Z}^{\lambda}_{0,i}\right)=\left(1,-1\right)$ only if
$X_{0,i}+X_{1,i}<0,\quad\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<\frac{\lambda\sigma^{2}}{2\mu}\left(1+\delta_{n}\right),\quad X_{1,i}<-\frac{\lambda\sigma^{2}}{2\mu},$
or
$X_{0,i}+X_{1,i}>0,\quad-\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<-\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right),\quad X_{1,i}>\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right).$
This leads to
$\mathbb{I}\left(\check{Z}^{\lambda}_{0,i}=1,\widehat{Z}^{\lambda}_{0,i}=-1,{\cal E}_{\sf par}\right)$
$\leq\mathbb{I}\left(X_{0,i}+X_{1,i}<0,\quad\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<\frac{\lambda\sigma^{2}}{2\mu}\left(1+\delta_{n}\right),\quad X_{1,i}<-\frac{\lambda\sigma^{2}}{2\mu}\right)$
$+\mathbb{I}\left(X_{0,i}+X_{1,i}>0,\quad-\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<-\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right),\quad X_{1,i}>\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right)\right).$ (40)
We first consider the case when $(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,1\right)$. By Lemma 6 we have the following bound for the first term in (D.7):
$\mathbb{P}\left(X_{0,i}+X_{1,i}<0,\quad\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<\frac{\lambda\sigma^{2}}{2\mu}\left(1+\delta_{n}\right),\quad X_{1,i}<-\frac{\lambda\sigma^{2}}{2\mu}\mid(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,1\right)\right)$
$\lesssim\delta_{n}\left(\frac{\lambda}{\mu/\sigma}\right)\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1-\frac{\lambda}{2\mu^{2}/\sigma^{2}}\right)^{2}\right)\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right).$
For the second term in (D.7), by Lemma 6 we have
$\mathbb{P}\left(X_{0,i}+X_{1,i}>0,-\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<-\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right),X_{1,i}>\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right)\mid(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,1\right)\right)$
$\lesssim\delta^{2}_{n}\left(\frac{\lambda}{\mu/\sigma}\right)^{2}\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1+\frac{\lambda\sigma^{2}}{2\mu^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right)\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1-\frac{\lambda\sigma^{2}}{2\mu^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right).$
Using Lemma 5 and $\delta_{n}=o\left(\left(\mu/\sigma\right)^{-2}\right)$, we arrive at
$\mathbb{P}\left(\check{Z}^{\lambda}_{0,i}=1,\widehat{Z}^{\lambda}_{0,i}=-1\mid(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,1\right)\right)\lesssim\frac{\delta_{n}\lambda\left(1+\delta_{n}\lambda\right)}{\mu/\sigma}\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right).$
When $(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,-1\right)$, for the first term in (D.7) by Lemma 6 we have
$\mathbb{P}\left(X_{0,i}+X_{1,i}<0,\quad\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<\frac{\lambda\sigma^{2}}{2\mu}\left(1+\delta_{n}\right),\quad X_{1,i}<-\frac{\lambda\sigma^{2}}{2\mu}\mid(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,-1\right)\right)$
$\lesssim\delta_{n}\left(\frac{\lambda}{\mu/\sigma}\right)\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1-\frac{\lambda}{2\mu^{2}/\sigma^{2}}\right)^{2}\right)\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right).$
For the second term in (D.7) by Lemma 6 we have
$\mathbb{P}\left(X_{0,i}+X_{1,i}>0,-\frac{\lambda\sigma^{2}}{2\mu}<X_{0,i}<-\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right),X_{1,i}>\frac{\lambda\sigma^{2}}{2\mu}\left(1-\delta_{n}\right)\mid(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,-1\right)\right)$
$\lesssim\delta^{2}_{n}\left(\frac{\lambda}{\mu/\sigma}\right)^{2}\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1+\frac{\lambda\sigma^{2}}{2\mu^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right)\exp\left(-\frac{\mu^{2}}{2\sigma^{2}}\left(1+\frac{\lambda\sigma^{2}}{2\mu^{2}}\right)^{2}\left(1-2\delta_{n}\right)\right).$
Using Lemma 5 and $\delta_{n}=o\left(\left(\mu/\sigma\right)^{-2}\right)$, we arrive at
$\mathbb{P}\left(\check{Z}^{\lambda}_{0,i}=1,\widehat{Z}^{\lambda}_{0,i}=-1\mid(Z_{0,i}^{*},Z_{1,i}^{*})=\left(1,-1\right),{\cal E}_{\sf par}\right)\lesssim\delta_{n}\left(\frac{\lambda}{\mu/\sigma}\right)\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right).$
The case for $(Z_{0,i}^{*},Z_{1,i}^{*})=\left(-1,1\right)$ and $(Z_{0,i}^{*},Z_{1,i}^{*})=\left(-1,-1\right)$ are almost the same and hence omitted. We finally reach that
$\mathbb{P}\left(\check{Z}^{\lambda}_{0,i}=1,\widehat{Z}^{\lambda}_{0,i}=-1,{\cal E}_{\sf par}\right)\lesssim\frac{\delta_{n}\lambda\left(1+\delta_{n}\lambda\right)}{\mu/\sigma}\left[\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)+\varepsilon\cdot\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right)\right]$
Similarly, we can also the same bound for $\mathbb{P}\left(\check{Z}^{\lambda}_{0,i}=1,\widehat{Z}^{\lambda}_{0,i}=-1\right)$ and conclude that
$\mathbb{P}\left(\widehat{Z}^{\lambda}_{0,i}\neq\check{Z}^{\lambda}_{0,i}\right)\leq\frac{C\delta_{n}\lambda\left(1+\delta_{n}\lambda\right)}{\mu/\sigma}\left[\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)+\varepsilon\cdot\Phi\left(-\frac{\mu}{\sigma}+\frac{\lambda}{2\mu/\sigma}\right)\right]+n^{-C_{\sf par}}.$
D.8 Proof of Lemma 10
Notice that for $i\in[n]$ we can write
$\begin{cases}X_{0,i}=\mu Z^{*}_{0,i}+\sigma V_{0,i}\\ \bar{X}_{1,i}=\mu Z^{*}_{0,i}+\sigma V_{1,i}\end{cases}\quad\begin{cases}\widetilde{X}_{0,i}=\widehat{\mu}Z^{*}_{0,i}+\widehat{\sigma}V_{0,i}\\ \widetilde{X}_{1,i}=\widehat{\mu}Z^{*}_{0,i}+\widehat{\sigma}V_{1,i}\end{cases}.$
[…]
We first consider the case when $\lambda=\infty$. For $\forall i\in[n]$, notice that
$\bar{Z}^{\infty}_{0,i}$ $=\operatorname*{arg\,min}_{u\in\{\pm 1\}}\left\{-\frac{\mu}{\sigma^{2}}u\left(X_{0,i}+\bar{X}_{1,i}\right)\right\}=\operatorname*{arg\,min}_{u\in\{\pm 1\}}\left\{-2\left(\frac{\mu}{\sigma}\right)^{2}uZ^{*}_{0,i}-\frac{\mu}{\sigma}u\left(V_{0,i}+V_{1,i}\right)\right\}$
$=\text{sign}\left(\frac{2\mu}{\sigma}Z^{*}_{0,i}+V_{0,i}+V_{1,i}\right)=Z^{*}_{0,i}\cdot\text{sign}\left(1+\frac{\sigma}{\sqrt{2}\mu}\frac{\left(V_{0,i}+V_{1,i}\right)Z^{*}_{0,i}}{\sqrt{2}}\right).$
Similarly, we have
$\widetilde{Z}^{\infty}_{0,i}$ $=Z^{*}_{0,i}\cdot\text{sign}\left(1+\frac{\widehat{\sigma}}{\sqrt{2}\widehat{\mu}}\frac{\left(V_{0,i}+V_{1,i}\right)Z^{*}_{0,i}}{\sqrt{2}}\right).$
Denote $\eta:=\frac{\left(V_{0,i}+V_{1,i}\right)Z^{*}_{0,i}}{\sqrt{2}}\sim N(0,1)$, then we have $\bar{Z}^{\infty}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}$ if the following event occurs:
$\left\{\eta>-\sqrt{2}\frac{\mu}{\sigma},\quad\eta<-\sqrt{2}\frac{\widehat{\mu}}{\widehat{\sigma}}\right\}\bigcup\left\{\eta<-\sqrt{2}\frac{\mu}{\sigma},\quad\eta>-\sqrt{2}\frac{\widehat{\mu}}{\widehat{\sigma}}\right\}.$
Hence we can arrive at
$\mathbb{I}\left(\bar{Z}^{\infty}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}\right)$ $\leq\mathbb{I}\left(\eta>-\sqrt{2}\frac{\mu}{\sigma},\quad\eta<-\sqrt{2}\frac{\widehat{\mu}}{\widehat{\sigma}}\right)+\mathbb{I}\left(\eta<-\sqrt{2}\frac{\mu}{\sigma},\quad\eta>-\sqrt{2}\frac{\widehat{\mu}}{\widehat{\sigma}}\right).$ (41)
It easy to obtain that (by modifying Lemma 7 and utilizing Lemma 5)
$\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq Z_{0,i}^{*}\right)\geq\Phi\left(-\frac{\sqrt{2}\mu}{\sigma}\right)\gtrsim\frac{1}{\mu/\sigma}\exp\left(-\frac{\mu^{2}}{\sigma^{2}}\right).$
On the other hand, Lemma 6, (41) and Assumption 4 imply that
$\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}\right)$
$\leq\mathbb{P}\left(-\sqrt{2}\frac{\mu}{\sigma}<\eta<-\sqrt{2}\frac{\mu}{\sigma}\left(1-\delta_{n}\right)\right)+\mathbb{P}\left(-\sqrt{2}\frac{\mu}{\sigma}\left(1+\delta_{n}\right)<\eta<-\sqrt{2}\frac{\mu}{\sigma}\right)+n^{-C_{\sf par}}$
$\leq\delta_{n}\frac{2\sqrt{2}\mu}{\sigma}\exp\left(-\frac{\mu^{2}}{\sigma^{2}}\left(1-2\delta_{n}\right)\right)+n^{-C_{\sf par}}$
Hence we can conclude that $\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq\widetilde{Z}^{\infty}_{0,i}\right)=o\left(\mathbb{P}\left(\bar{Z}^{\infty}_{0,i}\neq Z_{0,i}^{*}\right)\right)$ provided that $\delta_{n}=o\left(\left(\mu/\sigma\right)^{-2}\right)$. We then consider $\forall\lambda\in\Lambda$. For $\forall i\in[n]$, we have
$\left(\bar{Z}^{\lambda}_{0,i},\bar{Z}^{\lambda}_{1,i}\right)$ $=\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\frac{\mu}{\sigma^{2}}\left(uX_{0,i}+v\bar{X}_{1,i}\right)+\lambda\mathbb{I}\left(u\neq v\right)\right\}$
$=\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\left(\frac{\mu}{\sigma}\right)^{2}Z^{*}_{0,i}\left(u+v\right)-\frac{\mu}{\sigma}\left(uV_{0,i}+vV_{1,i}\right)+\lambda\mathbb{I}\left(u\neq v\right)\right\}.$
Hence $\bar{Z}^{\lambda}_{0,i}=1$ if the following event occurs:
$V_{1,i}>-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\quad V_{0,i}>-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\quad V_{0,i}+V_{1,i}>-\frac{2\mu}{\sigma}Z^{*}_{0,i},$
or
$V_{1,i}<-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\quad V_{0,i}>\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\quad V_{0,i}>V_{1,i}.$
On the other hand, we have $\widetilde{Z}^{\lambda}_{0,i}=-1$ if
$V_{1,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad V_{0,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad V_{0,i}+V_{1,i}<-\frac{2\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},$
or
$V_{1,i}>\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad V_{0,i}<-\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad V_{0,i}<V_{1,i}.$
Under (12), we have $\left(\bar{Z}^{\lambda}_{0,i},\widetilde{Z}^{\lambda}_{0,i}\right)=\left(1,-1\right)$ if
$-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{1,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},$
$-\frac{2\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}+V_{1,i}<-\frac{2\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},$
or
$V_{1,i}>\max\left\{-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i}\right\},\quad-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<-\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},$
$V_{0,i}+V_{1,i}>-\frac{2\mu}{\sigma}Z^{*}_{0,i},$
or
$V_{1,i}<\min\left\{-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i}\right\},\quad\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},$
$V_{0,i}+V_{1,i}<-\frac{2\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i}.$
This leads to
$\mathbb{I}\left(\bar{Z}^{\lambda}_{0,i}=1,\widetilde{Z}^{\lambda}_{0,i}=-1,{\cal E}_{\sf par}\right)$
$\leq\mathbb{I}\left(-\frac{2\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}+V_{1,i}<-\frac{2\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},~{}{\cal E}_{\sf par}\right)$
$+\mathbb{I}\left(V_{1,i}>\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<-\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},~{}{\cal E}_{\sf par}\right)$
$+\mathbb{I}\left(V_{1,i}<-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\quad\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},~{}{\cal E}_{\sf par}\right).$ (42)
By Lemma 7, we have
$\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}\right)$ $\geq\left[\Phi\left(-\frac{\mu}{\sigma}\right)\right]^{2}+\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)$
$\gtrsim\frac{1}{\mu^{2}/\sigma^{2}}\exp\left(-\frac{\mu^{2}}{\sigma^{2}}\right)+\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right).$ (43)
It suffices to verify that the expectation of first three terms in (D.8) is negligible compared to (D.8). For the first term in (D.8), by Lemma 6 we have
$\mathbb{P}\left(-\frac{2\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}+V_{1,i}<-\frac{2\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},~{}{\cal E}_{\sf par}\right)\lesssim\delta_{n}\left(\frac{\sqrt{2}\mu}{\sigma}\right)\exp\left(-\frac{\mu^{2}}{\sigma^{2}}\left(1-2\delta_{n}\right)\right)+n^{-C_{\sf par}},$
which is of negligible given $\delta_{n}=o\left(\left(\mu/\sigma\right)^{-3}\right)$. We first consider the case when $Z_{0,i}^{*}=1$. For the second term in (D.8), we have
$\mathbb{P}\left(V_{1,i}>\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<-\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},~{}{\cal E}_{\sf par}\right)$
$\lesssim\delta_{n}\left(\frac{\lambda\sigma}{2\mu}+\frac{\mu}{\sigma}\right)\exp\left(-\frac{1}{2}\left(\frac{\mu}{\sigma}+\frac{\lambda\sigma}{2\mu}\right)^{2}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}+\delta_{n}\left(\frac{\lambda\sigma}{2\mu}+\frac{\mu}{\sigma}\right)\right)+n^{-C_{\sf par}}.$ (44)
When $0<\lambda<2\left(\frac{\mu}{\sigma}\right)^{2}\frac{1+\delta_{n}}{1-\delta_{n}}$, we have $\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)\geq\Phi\left(-\frac{2\delta_{n}}{1-\delta_{n}}\frac{\mu}{\sigma}\right)\gtrsim 1$ and $\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}+\delta_{n}\left(\frac{\lambda\sigma}{2\mu}+\frac{\mu}{\sigma}\right)\right)\leq 1$. When $\lambda>2\left(\frac{\mu}{\sigma}\right)^{2}\frac{1+\delta_{n}}{1-\delta_{n}}$, by Lemma 5 and 6 we have
$\frac{\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}+\delta_{n}\left(\frac{\lambda\sigma}{2\mu}+\frac{\mu}{\sigma}\right)\right)}{\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)}\lesssim\delta_{n}\left(\frac{\lambda}{2\mu/\sigma}\right)^{2}.$
Combining two cases, we conclude that (D.8) negligible compared to (D.8) given $r<2$, $\delta_{n}=o\left(\left(\frac{\mu/\sigma}{\log n}\right)^{2}\right)$ and $\lambda\lesssim\log n$. For the third term in (D.8), we have
$\mathbb{P}\left(V_{1,i}<-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\quad\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},~{}{\cal E}_{\sf par}\right)$
$\leq\delta_{n}\left(\frac{\lambda\sigma}{2\mu}+\frac{\mu}{\sigma}\right)\exp\left(-\frac{1}{2}\left(\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)^{2}\right)\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)+n^{-C_{\sf par}}.$
which is of negligible given $\delta_{n}=o\left(\left(\mu/\sigma\right)^{-2}\right)$ by Lemma 5. We then consider $Z_{0,i}^{*}=-1$. For the second term in (D.8), we have
$\mathbb{P}\left(V_{1,i}>\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},\quad-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<-\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},~{}{\cal E}_{\sf par}\right)$
$\lesssim\delta_{n}\left(\frac{\lambda\sigma}{2\mu}+\frac{\mu}{\sigma}\right)\exp\left(-\frac{1}{2}\left(\frac{\mu}{\sigma}-\frac{\lambda\sigma}{2\mu}\right)^{2}\right)\Phi\left(-\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}+\delta_{n}\left(\frac{\lambda\sigma}{2\mu}+\frac{\mu}{\sigma}\right)\right)+n^{-C_{\sf par}},$
which is of negligible compared to (D.8) given $r<2$, $\delta_{n}=o\left(\left(\frac{\mu/\sigma}{\log n}\right)^{2}\right)$ and $\lambda\lesssim\log n$ by a similar argument to (D.8). For the last term in (D.8), we have
$\mathbb{P}\left(V_{1,i}<-\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i},\quad\frac{\lambda\sigma}{2\mu}-\frac{\mu}{\sigma}Z^{*}_{0,i}<V_{0,i}<\frac{\lambda\widehat{\sigma}}{2\widehat{\mu}}-\frac{\widehat{\mu}}{\widehat{\sigma}}Z^{*}_{0,i},~{}{\cal E}_{\sf par}\right)$
$\lesssim\delta_{n}\left(\frac{\lambda\sigma}{2\mu}+\frac{\mu}{\sigma}\right)\exp\left(-\frac{1}{2}\left(\frac{\mu}{\sigma}+\frac{\lambda\sigma}{2\mu}\right)^{2}\right)\Phi\left(\frac{\mu}{\sigma}-\frac{\lambda}{2\mu/\sigma}\right)+n^{-C_{\sf par}}.$
which is of negligible compared to (D.8) provided that $r<2$, $\delta_{n}=o\left(\left(\frac{\mu/\sigma}{\log n}\right)^{2}\right)$ and $\lambda\lesssim\log n$. Collecting the cases for $Z_{0,i}^{*}=1$ and $Z_{0,i}^{*}=-1$, we arrive at
$\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}=1,\widetilde{Z}^{\lambda}_{0,i}=-1\right)=o\left(\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}\neq Z^{*}_{0,i}\right)\right).$
The case $\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}=-1,\widetilde{Z}^{\lambda}_{0,i}=1\right)$ is almost the same (due to symmetry) and hence omitted, we thereby complete the proof by noticing
$\mathbb{P}\left(\widetilde{Z}^{\lambda}_{0,i}\neq\bar{Z}^{\lambda}_{0,i}\right)=\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}=1,\widetilde{Z}^{\lambda}_{0,i}=-1\right)+\mathbb{P}\left(\bar{Z}^{\lambda}_{0,i}=-1,\widetilde{Z}^{\lambda}_{0,i}=1\right).$
D.9 Proof of Lemma 11
Without loss of generality, we assume $\widetilde{\mu}>0$, i.e., $\widehat{\bm{\mu}}^{\top}{\bm{\mu}}\geq 0$. By definition, we have
$\frac{\left|\left\|\widehat{\bm{\mu}}\right\|/{\widehat{\sigma}}-\mu/{\sigma}\right|}{\mu/\sigma}$ $=\left|\frac{\left\|\widehat{\bm{\mu}}\right\|^{2}}{\widehat{\bm{\mu}}^{\top}{\bm{\mu}}}\cdot\frac{\sigma}{\widehat{\sigma}}-1\right|\leq\left|\frac{\left\|\widehat{\bm{\mu}}\right\|^{2}}{\widehat{\bm{\mu}}^{\top}{\bm{\mu}}}-1\right|+\left|\frac{\sigma}{\widehat{\sigma}}-1\right|\leq\delta_{n},$
where $\left|{\sigma}/{\widehat{\sigma}}-1\right|\leq\delta_{n}/2$ by 3, and
$\left|\frac{\left\|\widehat{\bm{\mu}}\right\|^{2}}{{\widehat{\bm{\mu}}^{\top}{\bm{\mu}}}}-1\right|=\left|\frac{\left\|\widehat{\bm{\mu}}\right\|^{2}-{\widehat{\bm{\mu}}^{\top}{\bm{\mu}}}}{\widehat{\bm{\mu}}^{\top}{\bm{\mu}}}\right|\leq\frac{\left\|\widehat{\bm{\mu}}\right\|\left\|\widehat{\bm{\mu}}-{\bm{\mu}}\right\|}{\left\|{\bm{\mu}}\right\|^{2}-\left\|\widehat{\bm{\mu}}-{\bm{\mu}}\right\|\left\|{\bm{\mu}}\right\|}\leq\frac{\left(1+\delta_{n}/4\right)\delta_{n}/4\left\|{\bm{\mu}}\right\|^{2}}{\left(1-\delta_{n}/4\right)\left\|{\bm{\mu}}\right\|^{2}}\leq\frac{\delta_{n}}{2}.$
Similarly, we could obtain the bound for the other term.
D.10 Proof of Lemma 12
Let $\upsilon:=\vartheta/\left(\vartheta+2\right)$. By Lemma 4, with probability at least $1-\delta$ we have
$\frac{1}{n}\sum_{i=1}^{n}X_{0,i}\leq\left(1+\upsilon\right)p+\frac{7\log\left(2/\delta\right)}{6\upsilon n}\leq\left(1+\upsilon\right)\vartheta_{1}q+2\vartheta_{2}+\frac{7\log\left(2/\delta\right)}{6\upsilon n},$
$\frac{1}{n}\sum_{i=1}^{n}X_{1,i}\geq\left(1-\upsilon\right)q-\frac{7\log\left(2/\delta\right)}{6\upsilon n}.$
Hence with probability at least $1-\delta$ we have
$\frac{1}{n}\sum_{i=1}^{n}X_{0,i}$ $\leq\frac{1+\upsilon}{1-\upsilon}\vartheta_{1}\cdot\frac{1}{n}\sum_{i=1}^{n}X_{1,i}+\left(1+\frac{1+\upsilon}{1-\upsilon}\vartheta_{1}\right)\frac{7\log\left(2/\delta\right)}{6\upsilon n}+2\vartheta_{2}$
$=\left(1+\vartheta\right)\vartheta_{1}\cdot\frac{1}{n}\sum_{i=1}^{n}X_{1,i}+\left(1+\frac{2}{\vartheta}\right)\left(1+\left(1+\vartheta\right)\vartheta_{1}\right)\frac{7\log\left(2/\delta\right)}{6n}+2\vartheta_{2}$
$\leq\frac{\left(1+\vartheta\right)\vartheta_{1}}{n}\sum_{i=1}^{n}X_{1,i}+\frac{7\log\left(2/\delta\right)}{2\vartheta n}\left(1+2\vartheta_{1}\right)+2\vartheta_{2},$
where we’ve used $\vartheta=\frac{2\upsilon}{1-\upsilon}$ in the second equality and $\vartheta<1$ in the last inequality.
D.11 Proof of Lemma 13
We first show the upper bound for $Q_{1-{\zeta}/{2}}(\widehat{F}_{N})$. By definition, we have
$\mathbb{P}\left(X\leq Q_{1-\zeta/3}\left(F\right)\right)\geq 1-\zeta/3.$
By Hoeffding’s inequality, we obtain that with probability at least $21-\eta/$ 2,
$\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(X_{i}\leq Q_{1-\zeta}\left(F\right)\right)\geq 1-\zeta/3-C\sqrt{\frac{\log\left(1/\eta\right)}{N}}\geq 1-\zeta/2,$
provided that $N\gtrsim\zeta^{-2}\log\left(1/\eta\right)$. On that event, by definition of $Q_{1-\zeta/2}(\widehat{F}_{N})$, we have $Q_{1-\zeta/2}(\widehat{F}_{N})\leq Q_{1-\zeta/3}\left(F\right)$.
It remains to show the lower bound. By definition of $Q_{1-\zeta}(F)$, we have that
$\mathbb{P}\left(X_{i}<Q_{1-\zeta}\left(F\right)\right)\leq 1-\zeta.$
Hoeffding’s inequality tells us that with probability at least $21-\eta/$ 2,
$\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(X_{i}<Q_{1-\zeta}\left(F\right)\right)\leq\mathbb{P}\left(X<Q_{1-\zeta}\left(F\right)\right)+C\sqrt{\frac{\log\left(1/\eta\right)}{N}}\leq 1-2\zeta/3<1-\zeta/2$
provided that $N\gtrsim\zeta^{-2}\log\left(1/\eta\right)$. Hence on that event,
  • Less than $21-\zeta/$ 2 fraction of $X_{i}$’s belongs to $(-\infty,Q_{1-\zeta}(F))$.
  • At least $21-\zeta/$ 2 fraction of $X_{i}$’s belongs to $(-\infty,Q_{1-\zeta/2}(\widehat{F}_{N})]$.
  • For any $\epsilon>0$, less than $21-\zeta/$ 2 fraction of $X_{i}$’s belongs to $(-\infty,Q_{1-\zeta/2}(\widehat{F}_{N})-\epsilon]$
We can conclude that $Q_{1-\zeta}(F)\leq Q_{1-\zeta/2}(\widehat{F}_{N})$.